### v1

# MSCCL++: Rethinking GPU Communication Abstractions for Cutting-edge AI Applications

Abstract
Modern cutting-edge AI applications are being developed over fast-evolving, heterogeneous, nascent hardware devices. This requires frequent reworking of the AI software stack to adopt bottom-up changes from new hardware, which takes time for general-purpose software libraries. Consequently, real applications often develop custom software stacks optimized for their specific workloads and hardware. Custom stacks help quick development and optimization, but incur a lot of redundant efforts across applications in writing non-portable code. This paper discusses an alternative communication library interface for AI applications that offers both portability and performance by reducing redundant efforts while maintaining flexibility for customization. We present MSCCL++, a novel abstraction of GPU communication based on separation of concerns: (1) a primitive interface provides a minimal hardware abstraction as a common ground for software and hardware developers to write custom communication, and (2) higher-level portable interfaces and specialized implementations enable optimization for different hardware environments. This approach makes the primitive interface reusable across applications while enabling highly flexible optimization. Compared to state-of-the-art baselines (NCCL, RCCL, and MSCCL), MSCCL++ achieves speedups of up to 3.8$\times$ for collective communication and up to 15% for real-world AI inference workloads. MSCCL++ is in production of multiple AI services provided by Microsoft Azure, and is also adopted by RCCL, the GPU collective communication library maintained by AMD. MSCCL++ is open-source and available at https://github.com/microsoft/mscclpp.
Now at xAI. Now at Microsoft AI.
1 Introduction
GPU communication has become a key area of optimization for high-performance AI applications. For instance, modern Large Language Models (LLMs) [11, 5] are often designed to run a single inference task across tens of GPUs (spanning a few physical nodes) to distribute computation and minimize the end-to-end latency. Training tasks often scale much further up to tens of thousands of GPUs [11]. GPUs are connected and communicate with each other through either intra-node (PCIe, NVLink [24], xGMI [3], etc.) or inter-node (Ethernet or InfiniBand (IB) [1]) links. Efficient communication is challenging due to the collective communication patterns of AI workloads that frequently distributes and gathers data across all GPUs simultaneously. For instance, communication patterns like AllReduce, AllGather, AllToAll, Broadcast, and ReduceScatter [38] are frequently used in modern AI applications to distribute computing workload over many GPUs and collect computed results into a single GPU.
A substantial portion of the end-to-end latency is often attributed to GPU communication in real workloads, usually in the 10% - 40% range of end-to-end LLM workloads. For example, Mixture-of-Experts [14, 12] layers with experts distributed across 16 GPUs over two nodes may spend 40% of end-to-end training time for AllToAll communication. Similarly, a GPT-3 [5] model inference may spend 30% of the end-to-end time for AllReduce communication. A vast amount of efforts has been made to optimize GPU communication, such as developing efficient communication algorithms (routing paths and transfer scheduling) [32, 7, 18], overlapping communication with computation [15, 37], and other lower-level optimizations in the stack [21, 4].
Despite these efforts, achieving high communication performance is still challenging and time-consuming in real applications. Specifically, practitioners still need to write custom communication code to achieve the best performance, often from scratch. For instance, TensorRT-LLM [25], a popular LLM framework featured by NVIDIA in recent years, implements custom AllReduce communication methods from scratch. This is notable given that there already exists a popular GPU communication library developed by NVIDIA for many years, the NVIDIA Collective Communication Library (NCCL) [21]. The custom communication of TensorRT-LLM outperforms NCCL in a wide range of LLM scenarios, especially when the data size is relatively small, while TensorRT-LLM still uses NCCL for larger data sizes. This raises a simple question: why should it be a new ad-hoc stack from scratch, instead of developing on top of an existing stack?
Surprisingly, such ad-hoc software development has already become a widely common practice in cutting-edge AI applications, not only in GPU communication but in general AI workloads. From our empirical studies, we find that the root cause lies in the fast evolving hardware. The enormous computing demand by modern AI applications is pushing the industry to aggressively upgrade chips and interconnects, which comes up with powerful but immature hardware that is dramatically different from the previous generation. Cutting-edge AI applications are usually targeted for deployment over the latest hardware to obtain the best performance and efficiency. Thus, it is not surprising for existing general-purpose libraries (such as NCCL) to perform only sub-optimal for those applications, because it takes time for those libraries to optimize the wide range of workload scenarios over diverse hardware environments. Consequently, production implementations go with custom quick-and-dirty optimization specific to their own workloads and hardware environments. This incurs redundant development efforts across applications and writes a lot of non-portable code, which leads to further development effort when a new hardware becomes available.
In this paper, we discuss alternative interface design for AI software libraries to support such custom optimization practices by reducing redundant efforts without harming the flexibility of customization. The key idea is to separate high-level abstractions and optimizations from primitive hardware abstractions. Specifically, we have two key observations on what software developers need in real practices. First, software developers usually work closely with hardware experts to design optimization strategies based on the latest hardware features. Therefore, we need a low-level interface that minimally abstracts the hardware, which provides a common ground where both software and hardware developers easily understand and co-work on. Second, most of the existing abstractions and optimizations made in the software stack are not useful or even hinder optimization in real practices. As developers spend most of their time on new hardware and workloads, we need a highly flexible interface where they can try various design and optimization strategies over, rather than sophisticated abstractions based on existing hardware and workloads.
Based on this insight, we propose MSCCL++, [^3] Microsoft Collective Communication Library ++, pronounced em-sickle plus-plus. a novel GPU communication stack designed for high-performance AI applications. Unlike existing libraries, MSCCL++ exposes the most fundamental building blocks of communication as a user interface, namely the primitive interface (MSCCL++ Primitive API). This interface is very close to GPU hardware as a shallow layer over low-level GPU instructions, hence quick adoption of latest hardware features with low efforts. The primitive interface abstracts straightforward concepts of communication such as put, get, signal, and wait. This helps translate low-level hardware behaviors into higher-level insights (e.g., bandwidth and latency) for optimizing communication algorithms. The primitive interface is designed to be very flexible to enable various optimizations that have been difficult to implement with existing stacks. It provides zero-copy, one-sided, and asynchronous communication abstractions for efficient communication both within and across nodes. These abstractions enable communication to be efficiently fused within compute kernels, allowing fine-grained overlapping between compute and communication.
While the primitive interface offers fine-grained optimization for GPU experts, MSCCL++ also provides higher-level interfaces over the primitive interface for quick and easy optimization. Inspired by MSCCLang [10], we design the MSCCL++ stack to separate the declaration of communication algorithms (MSCCL++ DSL (domain-specific language)) from the underlying implementation (MSCCL++ DSL Executor) using an easy-to-use interface (MSCCL++ DSL API). The MSCCL++ DSL accelerates developing custom communication algorithms in new hardware environments, while the performance is still on par with hand-written implementations in most cases. Lastly, for users with the least hardware expertise, we re-implement the popular NCCL API as-is over the MSCCL++ stack (MSCCL++ Collective API) so that applications can adopt it without changing the code at all. Although the power of MSCCL++ is fully realized with application-specific optimization that cannot be delivered with this API, we still observe significant performance benefits even for general collective communication operations.
2 Background and Motivation
2.1 Collective Communication for AI
Modern AI applications are usually based on large machine learning (ML) models that require training and serving on multiple accelerator devices (GPUs in this paper) in a collective manner. They distribute the model and data across multiple GPUs to reduce the computation and memory requirement per GPU. This distribution requires GPUs to share intermediate results at some points during the computation, which is done using collective communication operations. AllReduce is a common collective communication operation that sums up the partial results from all GPUs and broadcasts the computed result to all GPUs. AllReduce can be divided into two other collective operations: ReduceScatter and AllGather. ReduceScatter sums input buffers on all GPUs and distributes the output buffer equally on all GPUs. AllGather collects a distributed buffer from all GPUs and stores the full buffer on each GPU. ReduceScatter and AllGather are often used separately depending on the ML model architectures.
2.2 Limitations of Existing Abstractions
Existing collective communication libraries internally implement conventional networking abstractions such as send and recv inside GPU kernels. This would be a straightforward approach to generalize arbitrary communication algorithms in a hardware-agnostic manner. However, this would not be the best choice for performance, because such software abstractions are not well aligned with the GPU hardware abstractions. Instead, this paper proposes an API design from the perspective of GPU programming, which makes it more straightforward to efficiently utilize the GPU hardware.
To be specific, we provide an overview of the NCCL architecture in Section 2.2.1 and discuss its inefficiencies in Section 2.2.2. This would be a representative example of existing collective communication libraries for GPUs, because other popular libraries, such as ROCm Collective Communication Library (RCCL) [4] and Microsoft Collective Communication Library (MSCCL) [9], are designed based on NCCL and share the same limitations.
2.2.1 NCCL Architecture
NCCL provides a C++ API to setup connections between multiple GPUs and to run communication operations over GPUs. All participating processes (that usually hold one GPU each) of a distributed ML application initialize a NCCL handle to build connections beforehand, and then use the handle to call NCCL kernels (which perform collective operations on GPUs) during the execution. ML frameworks, such as PyTorch [27] and TensorFlow [2], wrap the NCCL API in Python, which are called by an ML model.
1global ringRS(in, nelem, ring)
2 send = ring.sendbuff; recv = ring.recvbuff;
3 sz = ring.buffSz;
4 prim = Primitives<half>
5 (tid, ring.buffSz, ring.prev, ring.next);
6
7 for (off = 0; off < nelem; off += sz)
8 //step 0: push data to the next GPU
9 idx = off + ring.ranks[ring.ranks-1]*sz;
10 prims.copy(in+idx, send, sz);
11 prims.send();
12
13 //k-2 steps: reduce and copy to next GPU
14 for (j = 2; j < ring.ranks; ++j)
15 rankDest = ring.ranks[ring.ranks-j];
16 idx = off + rankDest * buffSz;
17 prims.recv();
18 prim.reduce(in+idx, recv, send, sz, "+");
19 prims.send();
20
21 //step k-1: write the result for this rank
22 idx = off + ring.rank * buffSz;
23 prims.recv();
24 prims.reduce(recv, in+idx, sz, "+");
Figure 1: Ring ReduceScatter kernel in NCCL.
Figure 2: Workflow of the NCCL send and recv primitives between two GPUs connected using InfiniBand (IB). Even though IB NICs can directly read and write GPU memory, we need the CPU to initiate the transfer (ibv_post_send) and wait for the completion (ibv_poll_cq). GPU 0 waits before calling send for next batch of data, and GPU 1 waits before the receive buffer is ready.
NCCL Initialization. During initialization, NCCL obtains following information: (i) the rank of each process, (ii) number of processes, (iii) number of distributed nodes in the cluster, (iv) number of GPUs per node, (v) all links between GPUs in a node, including PCIe, NVLink, and xGMI, and (vi) all links between nodes, like InfiniBand (IB) and Ethernet. NCCL uses this information to create communication topologies such as ring and tree with bi-directional links. Then, NCCL allocates a send buffer that stores data to send to destination GPU and a receive buffer that stores data received from a source GPU. The metadata of the send and receive buffers on a GPU are shared with all peer GPUs, so that GPUs can either peer-to-peer access buffers of another GPU (through PCIe, NVLink, or xGMI) or use external hardware to copy data (through IB or Ethernet).
NCCL Kernel. NCCL implements a GPU kernel for all collective operations and communication algorithms (ring, tree, etc.). It chooses the best performing algorithm based on the data size of collective communication, namely the message size. NCCL kernels are built using four building block operations, called primitives, which are send, recv, copy, and reduce. (i) send copies data from the send buffer of source GPU to the receive buffer of destination GPU, (ii) recv waits until the transfer by source GPU’s send primitive has finished execution and unblocks send of the source GPU, (iii) copy copies source to destination buffer on the same GPU using GPU threads, and (iv) reduce does elementwise operation, like addition, on two buffers on the same GPU and write the result to another buffer. Subsequent calls to send blocks until destination GPUs have called recv to ensure that it is safe to write to send buffer. Additionally, NCCL defines extra operations that performs multiple of primitives in a single function call but that does not improve over the limitations we discuss in Section 2.2.2.
As an example, Figure 1 describes the NCCL kernel of a ring-based ReduceScatter for half precision floating point (FP16) using the NCCL primitives. [^4] The kernel has been simplified for readability. The kernel takes the input buffer, number of elements, a ring of GPUs, and writes output to the input buffer. The kernel obtains pointers to send and receive buffers, size of the buffer, and initializes a primitive object with information about both previous and next GPUs in the ring (lines 25). The main loop of the kernel performs ReduceScatter in a batch of send/receive buffer sizes (lines 724). The loop first copies each batch of input data and send it to the next GPU (lines 911). Then the loop goes through all other ranks in the ring and for each rank (i) computes the corresponding offset in the input buffer for the rank, (ii) receives the partially reduced data from the previous GPU in the receive buffer, (iii) reduces this received data with the input offset and stores in the send buffer, and (iv) sends this data to the next GPU in the rank (lines 1419). Finally, the GPU stores the data for offset on this GPU (line 24).
2.2.2 Limitations of NCCL Primitives
The send and recv primitives are not flexible enough to utilize a single-instruction multiple-thread (SIMT) architecture like GPUs efficiently. As GPU threads are executed in groups (warps in CUDA), [^5] Unlike CPU where each thread runs on a separate context and can be context-switched, GPU threads are grouped and executed altogether for a single instruction. Otherwise, it may underutilize the highly-parallel cores and the large memory bandwidth. GPU kernels usually case-by-case optimize the workload to carefully distribute tasks among threads and minimize synchronization overhead. Unfortunately, NCCL only implements a static abstraction that groups 128 - 640 threads to collectively call a single primitive and synchronize them at the end of the call. This static abstraction is not flexible enough to optimize the workload for different links and data sizes, and may introduce unnecessary overheads. We discuss the limitations of NCCL primitives below.
Wasted GPU cycles. The send and recv primitives block the calling threads until the data transfer is completed, and this may waste a bunch of GPU threads’ cycles in a busy-wait while loop. If the underlying link supports peer-to-peer memory access, send directly copies local data into the receive buffer of the destination GPU, which can efficiently utilize parallel GPU threads. However, if the link uses external hardware (e.g., NIC) for the transfer, send only wakes up a CPU thread (writes a few-byte flag) that initiates the transfer and busy-waits until the transfer is completed (see Figure 2). This is a light-weight task that only needs a single GPU thread, and it is a big waste of cycles to block many threads for this. An alternative would be making the send and recv asynchronous, which is also against the existing abstractions in NCCL.
Interconnect Optimizations. A link can allow different ways of transferring data. However, NCCL supports only one mode of data transfer for each link. For example, intra-node links, like PCIe and NVLink, have two modes of data transfer: (i) thread-copy, where multiple GPU threads read from source GPU memory and write to destination GPU memory, and (ii) DMA-copy, where the CPU initiates the DMA engine of the GPU to copy from source memory to destination memory using cudaMemcpyDeviceToDevice. Comparing to that of DMA-copy, thread-copy achieves lower latency, but often utilizes less bandwidth. For example, in our AllGather experiments over 8 NVIDIA A100 80G GPUs, thread-copy achieves only up to 227 GB/s NVLink bandwidth, while DMA-copy achieves 263 GB/s (+15.8%). However, NCCL uses only thread-copy unless it is infeasible by hardware. [^6] NCCL uses DMA copy only when peer-to-peer access between GPUs is not supported by the hardware configuration. Moreover, using DMA-copy frees GPU threads to other work, such as computations, which can lead to better overlapping of computation and communication depending on the application.
Inflexible Synchronization. NCCL primitives are self-synchronized to ensure data consistency. However, such synchronizations are often too conservative to realize correct communication semantics. Consider a loop where $N$ GPUs produce their own data, and each GPU consumes data from all other GPUs, and start all over again. When a single buffer is written by producers and read by consumers, two types of cross-GPU barriers are required: after each producer to ensure the following remote consumer sees consistent data, and after each consumer to prevent the following remote producer from overwriting what is being read by this consumer. To reduce synchronization overhead, we can use two buffers in rotation and only keep first the type of barriers which still ensures consistency and prevents overwriting, but is not possible with self-synchronized NCCL primitives.
Co-optimization with Computation. The most common practices in ML frameworks call separate GPU kernels for computation (e.g., cuBLAS or cuDNN) and collective communication (e.g., NCCL). However, many recent works [29, 15, 12, 17, 8, 39, 37] have shown performance gains by breaking the barrier and co-optimizing computation and communication operations. Unfortunately, coupling NCCL kernels with computation kernels is challenging because NCCL does not design the send and recv primitives as a programmable interface. Since the primitives are blocking and use a FIFO queueing mechanism, making them directly work with other parts of the application is not straightforward and necessitates a memory barrier in between. A lower-level GPU communication abstraction would be needed to enable the co-optimizations.
2.3 Limitations of Existing User Interfaces
NCCL API is designed to provide end-to-end implementations of collective communication operations. This would be what most applications expect from the library, but this is fundamentally limited to support diverse performance-sensitive applications. Like other GPU workloads do, collective communication over GPUs should be differently optimized for different scenarios, depending on the input size, GPU architecture, networking topology, and the application characteristics [32, 12]. It is very difficult for the library developers to provide the best performance for all cases in a timely manner.
MSCCL [9] is a previous work that tackles this issue by allowing developers to implement their own custom communication algorithms using a DSL. However, MSCCL is still based on the NCCL primitives and does not provide enough flexibility to fully utilize the hardware resources, as described in Section 2.2.2. The MSCCL++ Primitive API we propose in this paper is designed at a lower level that barely abstracts the GPU hardware, which enables optimization in the hardware’s perspective. This would provide a more flexible and efficient way to optimize GPU communication. Inspired by MSCCL, we also provide a DSL for software developers to easily develop communication algorithms.
3 MSCCL++ Overview
Figure 3: Overview of MSCCL++.
3.1 Hierarchical User Interfaces
MSCCL++ provides three user APIs at hierarchical levels of communication abstractions, as shown in Figure 3: Primitive, DSL, and Collective APIs. The three levels provide different trade-offs in terms of programming effort or expertise and execution latency or performance. The closer the interface is to the hardware, the higher the programmer’s control on execution, so higher the performance and programming effort.
MSCCL++ Primitive API (or MSCCL++ API). This replaces NCCL’s primitive API with a set of core communication functions that are called from the GPU kernels. This is paired with the bootstrapping API called from the host (CPU) side, which is used for connection setup between GPUs. The primitive interface is the base layer that is used to implement the other two higher-level interfaces. Programmers can also use the primitive interface directly to implement application-specific or hardware-specific optimizations in their own GPU kernels, including those kernels that also perform computation. As the primitive interface is our key contribution, we focus on its details throughout Section 3 and 4.
MSCCL++ DSL API. We reimplement the MSCCLang [10] DSL over MSCCL++, which allows to write custom collective communication algorithms in a high-level language. This interface is aimed at users who want to generate communication kernels that are optimized for their own workloads (application and hardware). We elaborate in Section 4.3.
MSCCL++ Collective API. We reimplement the NCCL API, including its bootstrapping API, as-is over MSCCL++. It is aimed at users with the least expertise; they can simply replace NCCL/RCCL with the MSCCL++ Collective library without changing their application code. This library has the same limitations as that of NCCL (Section 2.3); i.e., it may not provide the best algorithm for certain workloads. For better performance, users can install their own optimized algorithms written using the MSCCL++ DSL API into this library.
3.2 Primitive Communication Abstractions
3.2.1 Communication Channels
Since the mode(s) of data transfer supported differs between interconnects, MSCCL++ defines an abstract channel for each data transfer mode supported by the hardware and exposes communication primitives specific to that channel that can be invoked directly from inside a GPU kernel. MSCCL++ defines three communication channels [^7] For readers who are familiar with NCCL: MSCCL++ channel does not correspond to that in NCCL.: PortChannel, MemoryChannel, and SwitchChannel, which correspond to port-mapped I/O, memory-mapped I/O, and switch-mapped I/O respectively. A PortChannel uses interconnect ports to communicate between GPUs; i.e., a GPU can initiate the data transfer through a port to another GPU (note that ports are controlled by dedicated hardware for I/O, such as DMA engines on GPUs or RDMA NICs). A MemoryChannel uses peer-to-peer memory access to communicate between GPUs; i.e., a GPU can directly access another GPU’s memory. A SwitchChannel uses interconnection-switch-enabled multimem memory access to communicate between GPUs; i.e., a GPU can access the memory of multiple GPUs simultaneously through a switch.
The same interconnect can support multiple modes of data transfer, so multiple channels can be supported on the same interconnect; for example, all three channels are supported by NVLink, only PortChannel and MemoryChannel are supported by xGMI and PCIe, while only PortChannel is supported by InfiniBand. The implementation of each channel is specific to the interconnect. The SwitchChannel uses NVSwitch multimem instructions that support computation-on-switch. The MemoryChannel uses thread-copy data transfer mode over NVLink, xGMI, and PCIe. On the other hand, the PortChannel uses the DMA-copy data transfer mode over NVLink, xGMI, PCIe, and Infiniband. Current hardware interconnects require a CPU thread to initiate the data transfer. For example, transferring data over InfiniBand requires the CPU to call ibv_post_send that initiates the RDMA transfer from the sending GPU memory to the receiving GPU memory. However, such details are handled transparently by the implementation.
Generality. Port-mapped I/O, memory-mapped I/O, and switch-mapped I/O are complementary I/O methods in general computer architectures. PortChannel, MemoryChannel, and SwitchChannel are our abstractions to provide these I/O methods from inside GPU kernels. We believe that they are flexible enough to accommodate future hardware advances. For example, if future hardware interconnects supports initiating DMA-copy from within a GPU kernel, the same PortChannel API can target the new hardware.
1//async
2put(src0, dst1, size)
3//unsafe to reuse src0
4signal() //async
5flush() //sync
6//safe to reuse src0
(a) GPU-0
//unsafe to read dst1
wait() //sync
//safe to read dst1
(b) GPU-1
Figure 4: MSCCL++ data transfer abstractions. put asynchronously transfers data from one GPU to another. signal and wait synchronize data transfer between GPUs. flush ensures the completion of preceded data transfer.
3.2.2 Communication Primitives
All MSCCL++ primitives are defined as a method of a channel. The primitives are all conceptual and do not imply any specific implementation. In this section, we describe the primitives API for the PortChannel with an example. Primitives for the other channels, and the implementation for all the channels are explained in Section 4.
Borrowing the term from MPI [34], writing data to the peer’s side through a channel is called put. To provide our communication abstraction as close as possible to hardware capabilities, we design put to be zero-copy (i.e., no intermediate buffers), one-sided (i.e., initiated by a peer without participation of the other peer), and asynchronous. MSCCL++ provides other interfaces for synchronization purposes. Specifically, signal and wait provide a mechanism for synchronization (including memory consistency) across GPUs. flush is for local synchronization that ensures all previous put operations are already on the fly.
Figure 4 shows the semantics of four MSCCL++ primitives: put, signal, wait, and flush. The put primitive transfers data from src0 buffer of GPU-0 to dst1 buffer of GPU-1. GPU-0 calls a following signal primitive that signals GPU-1, which is asynchronous yet strictly ordered with the previous put. GPU-1 has to call the wait primitive before it can read dst1. GPU-0 calls the flush primitive to ensure that the previous put is on the fly and src0 can be safely reused.
MSCCL++ also supports a few fused primitives that can reduce the overhead of API calls. For example, since a put is usually followed by a signal, we provide the putWithSignal function that conducts both at once.
1global allPairsRS(count, gpus, channels[gpus])
2 sz = channel.scratchSz/gpus.num
3 count = count/gpus.num
4
5 for (off = 0; off < count; off += sz)
6 //Send 1/Nth data to each GPU
7 for (g = 0; g < gpus.num; g++)
8 idx = off + g * count;
9 channels[g].put(idx, sz*g, sz)
10 channels[g].signal()
11
12 //Reduce ever pair of GPU
13 for (g = 0; g < gpus.num-1; g++)
14 channels[g].wait()
15 reduce(in + off, channels[g].scratch)
16
17 //barrier on all gpus
18 multiDeviceBarrier();
Figure 5: All-pairs ReduceScatter kernel in MSCCL++. Channels are initialized with source as input and destination as scratch buffer.
Example. Figure 5 shows MSCCL++ implementation of the all-pairs ReduceScatter algorithm, which is not used in NCCL. In this algorithm, all GPUs send their data at $\frac{i}{N}$ offset to GPU $i$, the $i^{th}$ GPU does elementwise reduction of the received data, and stores the reduced data. To store the received data, the implementation allocates a scratch buffer per GPU. The kernel’s main loop performs reduction in batches of the scratch size (line 518). On each GPU, the loop first put $\frac{1}{N}$ of data to all other GPUs’ scratch buffer and then signals the completion of transfer (line 910). Then a GPU waits for the data to arrive from other GPUs and reduce it with its part of input data (line 1315). Finally, a barrier among all GPUs is required to ensure scratch is not overwritten for next main loop offset (line 1818). Section 5 shows that this algorithm works better than the ring algorithm for smaller data sizes while the ring algorithm works better for large data sizes.
3.3 Advantages
The MSCCL++ Primitive API itself does not introduce any novel optimization techniques – instead, it offers new straightforward abstractions that enable quick optimization of GPU communication. The key advantage comes from exposing the primitive functionalities as a user interface, which introduces three benefits. First, it allows users to easily customize GPU communication algorithms for their own infrastructure and workloads (see following paragraphs). This is valuable in practice because the library developers may not be aware of the specific workloads and hardware configurations of their users, hence the difficulties of timely support of efficient implementations. Second, the library developers can quickly deliver new hardware support by leaving the algorithm design to the users who can do it better (see Section 4.5.1). Third, since the primitive interface is close to the hardware, it is more flexible to adopt new hardware features.
We carefully optimized GPU communication in the MSCCL++ DSL Executor and Collective kernels by leveraging the features in the MSCCL++ Primitive API. Users of the Primitive API can also implement these optimizations directly. Section 5 shows how these optimizations help MSCCL++ obtains better performance than NCCL.
Asynchronous Communication. MSCCL++ contains separate primitives for data transfer and synchronization, thus, enabling asynchronous communication and batching the synchronization of multiple communication into a single synchronization. This abstraction has three benefits: (i) allows more algorithms to be specified than the synchronous communication in NCCL, (ii) enables faster collective communication performance by allowing an algorithm space that can batch several synchronizations into one, (iii) frees up resources of peer GPUs to do other meaningful work inside the application, and (iv) allows better inter-kernel optimizations.
Specialized GPU Kernels. By using channels for specific GPU interconnects, MSCCL++ allows users to specialize kernels for particular interconnects. Furthermore, MSCCL++ avoid extra copies in kernels, leading to a kernel with less code paths and zero register spills. For example, for eight A100 GPUs, the ring AllReduce of NCCL and MSCCL use 94 and 96 registers per thread, respectively, while an MSCCL++ version uses only 32 per thread. Due to these reasons, MSCCL++ kernels execute less instructions, avoid memory accesses due to register spills, and have better instruction cache hits.
Interconnect Optimizations. NCCL only allows a single mode for data transfer over NVLink at a time. In contrast, by using different channel types in MSCCL++ (PortChannel or MemoryChannel), user can utilize both thread-copy and DMA-copy data transfer mode over NVLink.
Computation-Communication Optimizations. Since MSCCL++ Primitive API is an in-kernel interface for communication, we can co-optimize both computation and the communication kernels, leading to better application performance. One such optimization is fusing communication with a GeMM kernel [15, 37, 17, 8, 39], so that the number of memory reads/writes decreases, and a GPU’s warp context switching enables overlapping of communication with the GeMM kernel. MSCCL++ Primitive API is an easy-to-use and efficient interface to implement this idea, and we remain the implementation as a future work.
4 Implementation
4.1 Initialization
In general, an MSCCL++ program runs as a multi-process on all nodes and GPUs within a node, such that each process is responsible for one GPU. Since MSCCL++ enables application-specific optimization, it provides CPU side interfaces for custom initialization of the collective communications used in an application. MSCCL++ provides a default initialization process that can be specialized by users for their application. The initialization process involves setting up connections between GPUs, finding the topology of GPU connections, and constructing channels and sharing scratch buffers between peer GPUs.
The first step in initialization is to create a communication channel between all distributed CPU processes to exchange metadata between processes. We call this communication channel a Bootstrap. A bootstrap consists of four virtual methods: (i) send to send data from one CPU process to another, (ii) recv to recv data from a CPU process, (iii) allGather to perform the AllGather collective communication, and (iv) barrier to synchronize all distributed processes. MSCCL++ implements the default bootstrap methods using POSIX sockets. If a user prefers another distributed communication protocol such as MPI or torch.distributed, it is straightforward to override bootstrap methods because all these methods directly correspond to existing distributed protocols.
The second step for application is to construct a communicator object using the bootstrap object. The application uses the communicator object to register buffers that are used to perform data transfers. These buffers can be either application buffers themselves or scratch buffers dedicated for communications. The communicator then creates relevant channel objects (PortChannel, MemoryChannel, or SwitchChannel) between GPUs based on physical links. For each GPU, these channel objects are initialized with buffers. Finally, these channels are shared among processes of each GPU.
4.2 Primitives for Channels and Protocols
A connection between GPUs in MSCCL++ is called a channel. PortChannel and MemoryChannel are peer-to-peer connections (i.e., between 2 GPUs), whereas SwitchChannel is a connection among a group/collection of (2 or more) GPUs. Figure 6 shows three types of channels and its interface in MSCCL++. A channel is initialized with the source and destination buffers, and an integer allocated on the GPU memory that servers as semaphore. In this section, we describe all three kinds of MSCCL++ channels.
1class Channel
2 protected:
3 //Src/Dst Buffers set during initialization
4 void* src, *dst;
5 //Semaphore alloced on this GPU
6 uint *semaphore; uint expectedVal;
7 public:
8 //Primitives
9 void signal() ; void wait()
10
11class PortChannel : public Channel
12 public:
13 void put(ulong dstOff, ulong srcOff, ulong sz,
14 uint tid, uint tids)
15
16template<Protocol protocol>
17class MemoryChannel : public Channel
18 public:
19 void put(ulong dstOff, ulong srcOff, ulong sz,
20 uint tid, uint tids, uint flag=0)
21 template<typename T>
22 T read(ulong off, uint tid, uint flag=0)
23
24 template<typename T>
25 void write(ulong off, uint tid, T elem,
26 uint flag=0)
27
28class SwitchChannel : public Channel
29 public:
30 void reduce(ulong dstOff, ulong srcOff,
31 ulong sz, uint tid, uint tids)
32 void broadcast(ulong dstOff, ulong srcOff,
33 ulong sz, uint tid, uint tids)
Figure 6: PortChannel, MemoryChannel, and SwitchChannel primitives in MSCCL++. Implementation of MemoryChannel primitives differ for the type of protocol: LL or HB.
4.2.1 PortChannel
A PortChannel implements primitives when data transfer is done over ports connected to GPU memory, such as using InfiniBand API’s ibv_post_send for IB or cudaMemcpy for DMA copy. Since data transfer over a port require the CPU to initiate the transfer, this channel creates a CPU thread for each GPU. The CPU thread reads data transfer or synchronization requests from a first-in-first-out request queue. The storage, head, and tail of a request queue are allocated using cudaMallocManaged, so that both CPU and GPU can access this data. By default we allocate a queue of 1024 requests but this size can be changed by the user. We now discuss the workflow of PortChannels using Figure 7.
Figure 7: PortChannel workflow for IB starting from ⓪ when GPU 0 calls put primitive to ⑧ when GPU 1 receives the data.
0
When the GPU calls a primitive, such as, put, then the first participating thread of the GPU pushes the request to the queue by writing at the head. Before writing to the queue, the GPU checks if the queue is filled, i.e., if the head value is more than the tail value. If the queue is filled then the GPU waits for the CPU to process atleast one request. 1 Then the GPU increments the head to the next element. 2 The CPU thread continously reads the element at the tail to see if there is a request from the GPU and when there is a request, then the CPU reads the request, zeros out the current element, and increments the tail. 3 Now the CPU thread will process the request. Below we explain the requests generated by primitives and how the CPU handle these requests.
Data Transfer. put pushes a put request in the queue. 4 The CPU processes this request by starting an RDMA transfer using ibv_post_send. Since this function is asynchronous, the CPU thread immediately returns. When the transfer is happening, peer-GPUs are free to execute code, thereby, improving the overall power efficiency of the system.
Synchronization. 5 signal pushes a signal request in the queue. 6 7 The CPU processes this request by atomically incrementing the semaphore on the receiving GPU using functions, like ibv_atomic_add for IB. The wait primitive on the receiving GPU do not create a request for its CPU, therefore, the receiving CPU is idle, rather wait continuously looks for the semaphore to reach a expected value in a while loop. 8 wait returns after the semaphore is incremented.
Flush. Primitive flush pushes a flush request in the queue and the first thread of the GPU wait until the queue head is equal to or more than the queue tail. The CPU process the flush request by waiting until all the previous data transfer and synchronization requests have been completed. For example, for IB we use ibv_poll_cq to get the status of all requests. After the flush request is complete, the GPU is free to re-write to the source buffer.
4.2.2 MemoryChannel
A MemoryChannel wraps data transfer methods that use thread-copy mechanism, i.e., directly use GPU threads for writing to peer GPUs memory.
Protocols. The channel provides two protocols, which define data transfer and synchronization technique to tune between low-latency and high-bandwidth: HB protocol provides a high-bandwidth but high-latency protocol, thus, is suitable for larger sizes, and LL protocol provides low-latency but low-bandwidth, thus, is suitable for smaller sizes. Both protocols achieves these properties by synchronizing on the suitable granularity of data.
(i) In HB protocol, peer GPUs transfer a large chunk of data and synchronize this chunk once using wait and signal primitives. Hence, the synchronization time is amortized over the transfer time leading to high bandwidth. However, since the receiving GPU needs to wait for the whole chunk before the GPU can process it, the HB protocol also has high latency.
(ii) In LL protocol, the receiving GPU synchronizes on fewer chunks of data being transferred and process the chunk as soon as it is received, thus, achieving lower latency. However, since the number of synchronization is proportional to the number of elements, this protocol provides lower-bandwidth. The LL protocol works as follows. The put primitive requires an extra integer flag value. For every $N-1$ elements written to the receiving GPU, put also write the flag. The receiving GPU uses the read primitive waits until the flag value at $N$ index of receiving buffer is set and then read and return the $N-1$ elements. However, we cannot arbitrarily use any $N$ because GPUs follow a weak memory consistency model where writes to different memory locations by multiple threads can be performed in any order. Therefore, we restrict $N$ to number of elements written by a single instruction, i.e., 4, 8, and 16 bytes memory accesses instructions. The flag value is decided based on the algorithm, such that, all flag values are distinct.
Implementation. A MemoryChannel is initialized with a protocol. A protocol defines the implementation of data transfer and synchronization primitives. We now describe the implementation of primitives based on protocol as follows:
Data Transfer. Primitive put reads data elements from source buffer and writes to destination buffer starting from different offsets. Similarly, the read primitive reads and return data from the destination buffer, and the write primitive writes data including the flag for the LL protocol to the destination buffer. To maximize bandwidth for the HB protocol, both primitives use 16-bytes loads and stores. For the LL protocol, both primitives also takes a flag and uses 8-byte load and stores by default or user supplied vector length. Since put both primitives are called by multiple threads (all or first few threads of the kernel), we achieve maximum bandwidth.
Synchronization. Synchronization primitives wait on a semaphore which is an integer allocated on the receiving GPU. The signal primitive atomically increments the semaphore of the receiving GPU and calls threadfence_system to ensure that writes by put and semaphore increments are made available in this order. The wait primitive performs a busy-wait while-loop that checks until the value of semaphore has reached the expected value. This wait is performed by the first thread of the kernel and all other threads wait on a kernel barrier. The channel tracks the expected value using its expectedValue member. Since after put returns memory writes are already in operation, src buffer can be reused. Therefore, the flush primitive is empty.
4.2.3 SwitchChannel
A SwitchChannel provides primitives for performing collective operations among GPUs. These operations usually require specialized hardware support. For example, NVIDIA NVLink 4.0 connects all H100 GPUs to a single NVSwitch and this NVSwitch can perform collective operations including reduce and broadcast. A SwitchChannel in MSCCL++ provides two primitives: reduce to add the corresponding elements of buffers residing on different GPUs and broadcast to send elements from a buffer on a GPU to all other GPUs. We now discuss the implementation of these primitives for NVIDIA NVSwitch, which provides in-network, switch-based aggregation and multicast capabilities using the NVLink SHARP (NVLS) technology [23]. However, we believe that these implementations can be generalized to other hardware in the future.
Reduce. This primitive takes a destination buffer as a local address on the local GPU and a source buffer allocated as a multimem address. A multimem address is a virtual address that points to different virtual addresses on each GPU that is a part of the channel/collective. The primitive goes through each element of the destination buffer on the local GPU and executes multimem.ld_reduce PTX instruction using the source element’s multimem address. The multimem instruction fetches the values from all the virtual memory addresses pointed by the multimem address to the switch, does the reduction on the switch, and returns the reduced value to the local GPU. The reduced output is obtained in a register and then written to the destination buffer.
Broadcast. This primitive takes a source buffer as a local address on the local GPU and a destination buffer as a multimem address. The primitive goes through each element of the source buffer on the local GPU, reads the element into a register, and executes the multimem.st PTX instruction using the source register and the destination element’s multimem address. The multimem instruction sends the register value to the switch, which broadcasts and stores the value to all the virtual memory addresses pointed by the multimem address.
4.3 DSL Implementation
MSCCL++ DSL API is a Python-based language that allows describing a communication algorithm at a high level. It converts the algorithm description into a sequence of instructions that can be executed by the DSL Executor, which is a GPU kernel that reads and runs instructions back-to-back. MSCCL++ DSL is an extension of the MSCCLang [10] that supports new instructions based on the MSCCL++ Primitive API and lifts a few restrictions of the language; e.g., we enable a single GPU thread block to access multiple GPUs at the same time, etc.
MSCCL++ DSL introduces a few limitations over the Primitive API. First, as the DSL introduces higher level instructions than the Primitive API, it may not be the most efficient way to implement. In our evaluation, DSL-written algorithms perform 3% worse than the hand-written ones on average, and is up to 18% worse in a corner case. Second, since a DSL-written algorithm can be run only by the provided executor kernel, it may not be straightforward to merge with other computation kernels if needed. Despite limitations, MSCCL++ DSL is still useful for quick prototyping of collective communication and easy-to-understand algorithm description.
4.4 Collective API and Algorithms
MSCCL++ Collective API is the highest level of API that MSCCL++ supports, and it is the same as the original NCCL API. It is implemented as a wrapper of the MSCCL++ DSL API, which means that users can install a DSL-written algorithm to be called by the NCCL API. Users can also use the default algorithms already provided by the API. At runtime, for each communication, MSCCL++ automatically selects the best performing algorithm.
To be specific, we provide high-level description of the AllReduce algorithms we implement using the MSCCL++ DSL API below. Each algorithm can be implemented in multiple ways for optimization depending on hardware and message size. For example, we implement eight different versions of the two-phase all-pairs (2PA) algorithm explained in a following paragraph, and we omit the details for brevity.
1. One-phase All-pairs (1PA). In all-pairs algorithms, all GPUs concurrently broadcast their own local data to all other GPUs. By one-phase, the algorithm sends all local data to all other GPUs, so that the reduction is done in a single phase. It is suitable for small message sizes where the synchronization overhead between GPUs is more critical than the redundant reduction and data traffic. As we use the 1PA algorithm only for very small message collectives within a single node, we implement only a single version of 1PA that uses the LL protocol with MemoryChannel. MSCCL++ can implement it much more efficient than baselines by removing unnecessary memory copies and synchronization as described in Section 2.2.2.
2. Two-phase All-pairs (2PA). By two-phase, the algorithm splits the AllReduce into two phases: the first for ReduceScatter (each of $N$ GPUs collect and reduce $1/N$ of the data) and the second for AllGather (each GPU broadcasts the reduced data to all other GPUs) [31]. Two-phase algorithms are more bandwidth-efficient and conducts less reduction than one-phase algorithms. In 2PA, the ReduceScatter and AllGather phases are done in the all-pairs manner each. We use 2PA for single-node collectives and implement multiple versions with combinations of options: MemoryChannel or PortChannel, LL or HB protocol, and put-based or get-based. Using MSCCL++, we can optimize 2PA in various ways that are not possible with baselines. For example, for up to a few MBs of messages, we exploit rotating buffers to reduce synchronization at the cost of using more memory space, as explained in Section 2.2.2. As another example, we can let a single thread group read data from multiple other GPUs at the same time. This allows efficient data reduction comparing to baselines that read data from different GPUs one-by-one, which syncrhonizes each reduction step.
3. Two-phase Hierarchical (2PH). Hierarchical algorithms exchange minimal data across nodes and do local collectives in each node to complete the operation. It can be faster than all-pairs by reducing the data traffic crossing the nodes. 2PH is a two-phase algorithm that performs ReduceScatter and AllGather in a hierarchical manner each. We use it for multi-node collectives and implement two versions. The first version is for small messages using LL protocol. Each node conducts local ReduceScatter that splits the data into the number of GPUs in a node. This requires to send more data across nodes and introduce redundant reduction, but it is faster for small messages by reducing synchronization steps. Cross-node communication is done in all-pairs manner. To utilize inter-/intra-node links at the same time, we pipeline the local collective with cross-node communication to overlap. The second version of 2PH is for large messages using HB protocol. Unlike the first version, to utilize the link bandwidth efficiently, the number of data chunks is the same as the number of GPUs. Similar to the first version, it performs all-pairs cross-node communication and local ReduceScatter in a pipelined manner.
4. Two-Phase All-Pairs Multimem (2PAM). 2PAM is a specialized algorithm for NVIDIA H100 GPUs that leverages the in-network, switch-based aggregation and multicast capabilities by the NVLink SHARP (NVLS) technology [23]. This algorithm simply walks through all data elements one-by-one in a loop to call two NVLS multimem instructions, one for in-network cumulating one element from all GPUs, and another for broadcasting the cumulated element.
4.5 Cross-hardware Implementation
This section introduces our experience in developing MSCCL++ across different hardware platforms and how we reduce development efforts for future hardware.
4.5.1 Quick Support for New GPUs
Extending MSCCL++ to support new hardware features is much easier compared to doing so in existing communication libraries, because the primitive interface is directly exposed to the user. This design accelerates the overall development because (1) the library developers only need to add a shallow layer of abstraction over low-level code that features key functionalities, and (2) the algorithm developers can write clever optimizations and fine-tuning with precise control of the hardware efficiency.
We provide an example of supporting the NVIDIA H100 GPU in MSCCL++, which delivers the hardware-based collective operations using NVSwitch (explained in Section 4.2.3). Since this feature introduced a very different concept from previous communication interfaces (i.e., PortChannel or MemoryChannel), we developed the SwitchChannel interface to support this feature. The development took only 8 weeks for two developers, including learning the basic usage of this feature, abstracting the feature as a new type of channel (i.e., SwitchChannel), and finally developing a new AllReduce algorithm using SwitchChannel that outperforms NCCL/MSCCL by more than 2.2$\times$ in average for message sizes of 1KB - 1GB.
As another example, we have developed MSCCL++ to support the AMD MI300x GPU, while the previous version supported only NVIDIA GPUs. The development took only 7 weeks for one developer: 3 weeks for basic AMD GPU support and 4 weeks to develop new AllReduce algorithms that outperform RCCL/MSCCL for message sizes of 1KB - 1GB. It is especially remarkable that the AMD-specific code in MSCCL++ (excluding makefiles and algorithms) is only less than 10 lines of code, while RCCL, the hard-forked NCCL for AMD GPUs, is substantially diverged from NCCL. This is possible because the low-level API of AMD GPUs (i.e., HIP) is almost the same as that of NVIDIA GPUs (i.e., CUDA), and the MSCCL++ Primitive API is only a shallow abstraction on top of the low-level API. This significantly reduces the effort to maintain and develop MSCCL++.
4.5.2 Managing Memory Consistency
Besides the proposed abstractions, MSCCL++ also puts particular considerations on its low-level implementation to reduce dependencies on certain hardware, thus reducing technical debt. As an example, we introduce our approach toward managing memory consistency. Ensuring memory consistency is critical as a communication library. In particular, fast-evolving chips and their compilers often consist of various bugs or undocumented behaviors that can lead to breaking memory consistency, which is often hard to detect and fix. This is a solid problem that prevents the quick adoption of the latest hardware techniques.
From our production efforts to tackle this problem, we found that the most effective and efficient way to ensure memory consistency is relying on the atomic operations, rather than leveraging other consistency models provided by hardware or compiler features. This makes it easier to blame the hardware or compiler when we find unexpected behaviors, [^8] Indeed, we found a bug of an atomic instruction in CUDA 12.x NVCC, which was reported to and fixed by NVIDIA. and also makes MSCCL++ more portable and maintainable. For example, we use atomic instructions that comply with the C++11 memory model instead of volatile instructions. As another example, for signal over RDMA, one may simply send a separate RDMA write request to send a flag to the target GPU, assuming that the underlying hardware and stacks are configured and working correctly to ensure ordered write. However, instead, MSCCL++ uses the atomic fetch-and-add operation of RDMA to ensure consistency from the software perspective (the impact on performance is negligible small).
Env. Name GPU Intra-node Link Network
A100-40G NVIDIA A100 (40G) (8x/node) NVLink 3.0 Mellanox HDR InfiniBand (200 Gb/s, 1x NIC/GPU)
A100-80G NVIDIA A100 (80G) (8x/node) NVLink 3.0 Mellanox HDR InfiniBand (200 Gb/s, 1x NIC/GPU)
H100 NVIDIA H100 (8x/node) NVLink 4.0 Quantum-2 CX7 InfiniBand (400 Gb/s, 1x NIC/GPU)
MI300x AMD MI300x (8x/node) Infinity Fabric Gen 4 Quantum-2 CX7 InfiniBand (400 Gb/s, 1x NIC/GPU)
Table 1: List of environments used for evaluation.
5 Evaluation
Environments. All presented numbers are collected from one of the environments listed in Table 1. Each node is equipped with 8 GPUs per node (either A100, H100, or MI300X), intra-node links between GPUs (either NVLink or Infinity Fabric (a.k.a. xGMI)), and inter-node InfiniBand links (one NIC per GPU). All NICs are connected to a single InfiniBand networking switch. NVIDIA GPUs use CUDA 12.4, while AMD GPUs use ROCm 6.2. For brevity, the rest of this paper refers to each environment by the name of GPU.
Baselines. We compare MSCCL++ with existing collective communication libraries, including NCCL 2.23.4 [21], RCCL 2.20.5 [4], [^9] As RCCL partly adopts MSCCL, to prevent confusion, we refer to MSCCL-disabled RCCL. and MSCCL 2.23 [19]. MSCCL borrows the stack implementation of NCCL/RCCL and supports custom algorithms. Therefore, MSCCL performs exactly the same as NCCL/RCCL if they use the same algorithm.
5.1 Collective Communication
We present collective communication performance of MSCCL++ and compare it with that of the baselines. We implement various GPU kernels based on the algorithms in Section 4.4, and we present the best number among all implementations for each message size as we do for NCCL, RCCL, and MSCCL as well. All NCCL, RCCL, and MSCCL numbers are fine-tuned for each environment and message size by adjusting their environment variables, such as the number of channels (affects the number of threads), chunk size (affects the size of data to be transferred at once), type of algorithm (such as ring, tree, or NVLS [23]), the topology (XML file that describes the intra-node link topology of GPUs), etc. For MSCCL, we use the fastest algorithm for each buffer size [20]. We leverage NCCL’s user buffer registration (i.e., ncclMemAlloc[26] and the CUDA/HIP Graph features for best performance.
For visibility, we separate the range of message sizes into small (up to 1MB, presented latency) and large (1MB and above, presented algorithm bandwidth (AlgoBW) [^10] Defined as the message size divided by the latency.) in the figures. The small message sizes represent inference scenarios (such as LLM token sampling a.k.a. decode [30]), while the large message sizes represent both training and inference scenarios (such as gradient accumulation during back-propagation, or LLM prompt processing, a.k.a. prefill [30]).
Figure 8: AllReduce, A100-40G, single-node (8 GPUs).
Figure 9: AllReduce, A100-40G, 2-node (16 GPUs).
Figure 10: AllReduce, A100-40G, 4-node (32 GPUs).
Figure 11: AllReduce, H100, single-node (8 GPUs).
Figure 12: AllReduce, MI300x, single-node (8 GPUs).
We compare the AllReduce performance over A100-40G nodes in Figure 8, 9, and 10. All data points are validated multiple times with separate trials. MSCCL++ outperforms baselines for both small messages (up to 3.5x and 2.1x faster over NCCL and MSCCL, respectively) and large messages (up to 1.6x and 1.4x faster over NCCL and MSCCL, respectively). MSCCL is faster than NCCL in most cases, and all the benefit comes from better algorithms. The gain is especially big for small message sizes where MSCCL uses all-pairs algorithms, while NCCL uses the ring algorithm that is worse in terms of latency. The gap between MSCCL and NCCL shrinks for very large message sizes where the performance is limited by the link bandwidth. MSCCL++ provides further benefits over MSCCL by implementing the algorithms in a more efficient way via the primitive API. All MSCCL and MSCCL++ numbers in the figures use the same collective algorithms and differ only by implementation in all cases. Even if MSCCL and MSCCL++ separately select the best-performing algorithm for each case, they come up with the same algorithm. Therefore, the gap between MSCCL and MSCCL++ directly shows the benefits of the proposed primitive API.
[…]
Figure 11 and 12 compare the single-node AllReduce performance over H100 and MI300x nodes. Similarly, MSCCL++ outperforms baselines by up to 3.8x and 2.2x for small and large message ranges, respectively. This shows that MSCCL++ is effective across different GPU architectures and links. The speedup over NCCL/MSCCL for large message sizes on H100 is especially interesting, because the algorithm used by MSCCL++ here is 2PAM (Section 4.4), which simply calls two NVLS multimem instructions element-wise (one for reduction and one for broadcast) in a loop. This shows that NCCL primitives are built with a big overhead of unnecessary components that can be avoided in MSCCL++.
5.2 LLM Inference Acceleration
Figure 13: Speedup for decodes with tensor parallelism = 8
[…]
6 Related Work
NVIDIA Collective Communications Library (NCCL) [21] and AMD RoCm Collective Communications Library (RCCL) [4] are vendor supplied libraries for NVIDIA and AMD GPUs respectively. Cowan et al. [9] improved over these libraries by allowing execution of collective algorithms specialized for a size and topology. In this paper, we showed that primitives provided by MSCCL++ performs better than the primitives of these libraries.
The primitive interface of MSCCL++ has similarities with some other works in terms of being a GPU-side API for communication. For example, the NVIDIA OpenSHMEM (NVSHMEM) [22] is a parallel work that provides primitive functions such as nvshmemx_putmem_warp as well as a few collective communication functions such as nvshmemx_broadcastmem_warp. However, we could not find any implementation where NVSHMEM outperforms NCCL (or MSCCL++) for collective communication, and qualitative comparison is difficult as NVSHMEM is not open source. ARK [13] is another work that proposes a GPU-side control plane for communication, but it is implemented as a monolithic end-to-end ML system rather than a standalone communication library.
UCX, an open-source framework developed by Shamis et al. [33] and libraries implementing the MPI standard [35] offer a comprehensive set of APIs for high-performance computing (HPC) environments, however, MSCCL++ is more specific to GPU acceleration.
Works like SCCL [6], TACCL [32], and TE-CCL [18] aim to accelerate GPU collective communication by synthesizing efficient data transfer algorithms for performing collective communication, while still using the vendor-provided primitives underneath. The primitives interface provided by MSCCL++ can be used with the generated algorithms to capture their benefit.
Finally, LLM inference frameworks like vLLM [16] and TensorRT-LLM [25] implement custom AllReduce kernels for achieving high performance communication. However, the custom implementations are not general-purpose and often only limited to single-node collective communication. MSCCL++, on the other hand, enables communication across multiple types of interconnects and has equivalent AllReduce performance to vLLM’s custom single-node AllReduce.
7 Conclusion
MSCCL++ is a novel GPU communication stack designed for high-performance AI applications. By exposing the primitive communication functionalities as straightforward user interfaces, MSCCL++ enable fine-grained optimizations for GPU experts, while also providing higher-level interfaces for quick optimizations. Moreover, such a design can reduce the overall development and optimization effort for GPU communication, and accelerates adoption of fast evolving hardware technologies. By implementing collective communication using the proposed MSCCL++ interfaces, we can achieve up to 3.8×3.8$\times$3.8 × speedup for standalone collectives, and up to $15\%$ speedup for end-to-end AI inference.

### v2

@@
 MSCCL++: Rethinking GPU Communication Abstractions for Cutting-Edge AI Applications
-Abstract
+Abstract.
 Modern cutting-edge AI applications are being developed over fast-evolving, heterogeneous, nascent hardware devices.
@@
 Consequently, real applications often develop custom software stacks optimized for their specific workloads and hardware.
-Custom stacks help quick development and optimization, but incur a lot of redundant efforts across applications in writing non-portable code.
+Custom stacks help in quick development and optimization, but incur a lot of redundant efforts across applications in writing non-portable code.
 This paper discusses an alternative communication library interface for AI applications that offers both portability and performance by reducing redundant efforts while maintaining flexibility for customization.
-We present MSCCL++, a novel abstraction of GPU communication based on separation of concerns: (1) a primitive interface provides a minimal hardware abstraction as a common ground for software and hardware developers to write custom communication, and (2) higher-level portable interfaces and specialized implementations enable optimization for different hardware environments.
+We present MSCCL++, a novel abstraction of GPU communication based on separation of concerns: (1) a primitive interface provides a minimal hardware abstraction as a common ground for software and hardware developers to write custom communication, and (2) higher-level portable interfaces and specialized implementations enable optimization for different workloads and hardware environments.
 This approach makes the primitive interface reusable across applications while enabling highly flexible optimization.
-Compared to state-of-the-art baselines (NCCL, RCCL, and MSCCL), MSCCL++ achieves speedups of up to 3.8$\times$ for collective communication and up to 15% for real-world AI inference workloads.
+Compared to state-of-the-art baselines (NCCL, RCCL, and MSCCL), MSCCL++ achieves speedups of up to 5.4$\times$ for collective communication and up to 15% for real-world AI inference workloads.
 MSCCL++ is in production of multiple AI services provided by Microsoft Azure, and is also adopted by RCCL, the GPU collective communication library maintained by AMD.
 MSCCL++ is open-source and available at https://github.com/microsoft/mscclpp.
-Now at xAI. Now at Microsoft AI.
-1 Introduction
+conference: ; ; Now at Microsoft AI. Now at xAI.
+1. Introduction
 GPU communication has become a key area of optimization for high-performance AI applications.
-For instance, modern Large Language Models (LLMs) [11, 5] are often designed to run a single inference task across tens of GPUs (spanning a few physical nodes) to distribute computation and minimize the end-to-end latency.
-Training tasks often scale much further up to tens of thousands of GPUs [11].
-GPUs are connected and communicate with each other through either intra-node (PCIe, NVLink [24], xGMI [3], etc.) or inter-node (Ethernet or InfiniBand (IB) [1]) links.
-Efficient communication is challenging due to the collective communication patterns of AI workloads that frequently distributes and gathers data across all GPUs simultaneously.
-For instance, communication patterns like AllReduce, AllGather, AllToAll, Broadcast, and ReduceScatter [38] are frequently used in modern AI applications to distribute computing workload over many GPUs and collect computed results into a single GPU.
+For instance, modern Large Language Models (LLMs) (Dubey et al., 2024; Brown et al., 2020) are often designed to run a single inference task on tens of GPUs (spanning a few physical nodes) to distribute computation and minimize end-to-end latency.
+Training tasks often scale much further up to tens of thousands of GPUs (Dubey et al., 2024).
+GPUs are connected and communicate with each other through either intra-node (PCIe, NVLink (NVIDIA, 2025d), xGMI (AMD, 2023), etc.) or inter-node (Ethernet or InfiniBand (IB) (inf, 2023)) links.
+Efficient communication is challenging due to the collective communication patterns of AI workloads that frequently distribute and gather data across all GPUs simultaneously.
+For example, communication patterns such as AllReduce, AllGather, AllToAll, Broadcast, and ReduceScatter (Wikipedia, 2025) are frequently used in modern AI applications to distribute computing workload over many GPUs and collect computed results into a single GPU.
 A substantial portion of the end-to-end latency is often attributed to GPU communication in real workloads, usually in the 10% - 40% range of end-to-end LLM workloads.
-For example, Mixture-of-Experts [14, 12] layers with experts distributed across 16 GPUs over two nodes may spend 40% of end-to-end training time for AllToAll communication.
-Similarly, a GPT-3 [5] model inference may spend 30% of the end-to-end time for AllReduce communication.
-A vast amount of efforts has been made to optimize GPU communication, such as developing efficient communication algorithms (routing paths and transfer scheduling) [32, 7, 18], overlapping communication with computation [15, 37], and other lower-level optimizations in the stack [21, 4].
-Despite these efforts, achieving high communication performance is still challenging and time-consuming in real applications.
+For example, a GPT-3 (Brown et al., 2020) model inference may spend 30% of the end-to-end time in AllReduce communication.
+Many efforts have been made to optimize GPU communication, such as developing efficient communication algorithms (routing paths and transfer scheduling) (Shah et al., 2023; Cao et al., 2024; Liu et al., 2024), overlapping communication with computation (Jangda et al., 2022; Wang et al., 2023), and other lower-level optimizations in the stack (NVIDIA, 2025a; AMD, 2025).
+Despite these efforts, achieving high communication performance is still challenging and time consuming in real applications.
 Specifically, practitioners still need to write custom communication code to achieve the best performance, often from scratch.
-For instance, TensorRT-LLM [25], a popular LLM framework featured by NVIDIA in recent years, implements custom AllReduce communication methods from scratch.
-This is notable given that there already exists a popular GPU communication library developed by NVIDIA for many years, the NVIDIA Collective Communication Library (NCCL) [21].
+For instance, TensorRT-LLM (NVIDIA, 2025e), a popular LLM framework featured by NVIDIA in recent years, implements custom AllReduce communication methods from scratch.
+This is notable given that there already exists a popular GPU communication library developed by NVIDIA for many years, the NVIDIA Collective Communication Library (NCCL) (NVIDIA, 2025a).
 The custom communication of TensorRT-LLM outperforms NCCL in a wide range of LLM scenarios, especially when the data size is relatively small, while TensorRT-LLM still uses NCCL for larger data sizes.
-This raises a simple question: why should it be a new ad-hoc stack from scratch, instead of developing on top of an existing stack?
-Surprisingly, such ad-hoc software development has already become a widely common practice in cutting-edge AI applications, not only in GPU communication but in general AI workloads.
+This raises a simple question: Why build a new ad hoc stack from scratch, instead of developing on top of an existing stack?
+Surprisingly, such ad hoc software development has already become a widely common practice in cutting-edge AI applications.
 From our empirical studies, we find that the root cause lies in the fast evolving hardware.
-The enormous computing demand by modern AI applications is pushing the industry to aggressively upgrade chips and interconnects, which comes up with powerful but immature hardware that is dramatically different from the previous generation.
+The enormous computing demand by modern AI applications is pushing the industry to aggressively upgrade chips and interconnects, resulting in powerful but immature hardware that is dramatically different from the previous generation.
 Cutting-edge AI applications are usually targeted for deployment over the latest hardware to obtain the best performance and efficiency.
-Thus, it is not surprising for existing general-purpose libraries (such as NCCL) to perform only sub-optimal for those applications, because it takes time for those libraries to optimize the wide range of workload scenarios over diverse hardware environments.
+Thus, it is not surprising for existing general-purpose libraries (such as NCCL) to perform sub-optimally for those applications because it takes time for those libraries to optimize for diverse workload scenarios.
 Consequently, production implementations go with custom quick-and-dirty optimization specific to their own workloads and hardware environments.
 This incurs redundant development efforts across applications and writes a lot of non-portable code, which leads to further development effort when a new hardware becomes available.
-In this paper, we discuss alternative interface design for AI software libraries to support such custom optimization practices by reducing redundant efforts without harming the flexibility of customization.
-The key idea is to separate high-level abstractions and optimizations from primitive hardware abstractions.
-Specifically, we have two key observations on what software developers need in real practices.
-First, software developers usually work closely with hardware experts to design optimization strategies based on the latest hardware features.
-Therefore, we need a low-level interface that minimally abstracts the hardware, which provides a common ground where both software and hardware developers easily understand and co-work on.
-Second, most of the existing abstractions and optimizations made in the software stack are not useful or even hinder optimization in real practices.
-As developers spend most of their time on new hardware and workloads, we need a highly flexible interface where they can try various design and optimization strategies over, rather than sophisticated abstractions based on existing hardware and workloads.
-Based on this insight, we propose MSCCL++, [^3] Microsoft Collective Communication Library ++, pronounced em-sickle plus-plus.
-a novel GPU communication stack designed for high-performance AI applications.
+In this paper, we discuss an alternative interface design for AI software libraries to support portability and custom optimizations by reducing redundant efforts without harming the flexibility of customization.
+The key idea is to separate high-level abstractions and optimizations from primitive hardware abstractions based on two observations.
+First, existing hardware abstractions like NCCL primitives try to hide hardware features to make it easier to program, but may hinder optimizations.
+Instead, a low-level, primitive interface that minimally abstracts the hardware would make it easier for: (1) application developers to optimize and fine-tune by enabling precise control of the hardware, and (2) library developers to support new hardware features by only providing a shallow layer of abstraction over it.
+Second, existing collective communication libraries (like NCCL) and DSLs (domain-specific languages; like MSCCLang (Cowan et al., 2023)) provide ease of use, but may lack workload-specific optimizations.
+Such higher level abstractions could provide portability and enable users to trade-off performance for programming effort.
+Based on these insights, we propose MSCCL++ (Microsoft Collective Communication Library ++, pronounced em-sickle plus-plus), a novel GPU communication stack designed for high-performance AI applications.
 Unlike existing libraries, MSCCL++ exposes the most fundamental building blocks of communication as a user interface, namely the primitive interface (MSCCL++ Primitive API).
-This interface is very close to GPU hardware as a shallow layer over low-level GPU instructions, hence quick adoption of latest hardware features with low efforts.
-The primitive interface abstracts straightforward concepts of communication such as put, get, signal, and wait.
-This helps translate low-level hardware behaviors into higher-level insights (e.g., bandwidth and latency) for optimizing communication algorithms.
+This interface is very close to GPU hardware as a shallow layer over low-level GPU instructions, thus enabling quick adoption of latest hardware features with low effort.
+The primitive interface abstracts straightforward concepts of communication such as put, signal, wait, and flush.
+This helps translate low-level hardware behaviors into higher-level insights (e.g., bandwidth and latency) for optimizing communication.
 The primitive interface is designed to be very flexible to enable various optimizations that have been difficult to implement with existing stacks.
-It provides zero-copy, one-sided, and asynchronous communication abstractions for efficient communication both within and across nodes.
-These abstractions enable communication to be efficiently fused within compute kernels, allowing fine-grained overlapping between compute and communication.
-While the primitive interface offers fine-grained optimization for GPU experts, MSCCL++ also provides higher-level interfaces over the primitive interface for quick and easy optimization.
-Inspired by MSCCLang [10], we design the MSCCL++ stack to separate the declaration of communication algorithms (MSCCL++ DSL (domain-specific language)) from the underlying implementation (MSCCL++ DSL Executor) using an easy-to-use interface (MSCCL++ DSL API).
-The MSCCL++ DSL accelerates developing custom communication algorithms in new hardware environments, while the performance is still on par with hand-written implementations in most cases.
-Lastly, for users with the least hardware expertise, we re-implement the popular NCCL API as-is over the MSCCL++ stack (MSCCL++ Collective API) so that applications can adopt it without changing the code at all.
-Although the power of MSCCL++ is fully realized with application-specific optimization that cannot be delivered with this API, we still observe significant performance benefits even for general collective communication operations.
-2 Background and Motivation
-2.1 Collective Communication for AI
+It provides zero-copy, one-sided, and asynchronous abstractions for efficient communication both within and across nodes.
+While the MSCCL++ Primitive API offers fine-grained optimization for GPU experts, MSCCL++ also provides higher-level interfaces that are easier to program.
+Inspired by MSCCLang (Cowan et al., 2023), we separate the declaration of communication algorithms, done using the MSCCL++ DSL API, from the underlying implementation of the algorithm, done by the MSCCL++ DSL Executor using MSCCL++ primitives.
+The MSCCL++ DSL accelerates the development of application-specific communication algorithms in new hardware environments.
+Lastly, for users with the least hardware expertise, we re-implement the popular NCCL API as-is (MSCCL++ Collective API) using MSCCL++ primitives so that applications can adopt it without changing the code at all.
+In our evaluation, we compare SpeCL with NCCL, ROCm Collective Communication Library (RCCL) (AMD, 2025) and Microsoft Collective Communication Library (MSCCL) (Cowan et al., 2022) on NVIDIA A100, NVIDIA H100, and AMD MI300x GPUs.
+We show that collective communication kernels like AllReduce and AllGather written using the MSCCL++ DSL API are up to 5.4×5.4$\times$5.4 × faster than the other libraries.
+We also re-implement some of these communication algorithms directly using the MSCCL++ Primitive API and show that those kernels are on average $3\%$ faster than their counterparts.
+These algorithm implementations are included in the MSCCL++ Collective library.
+These benefits translate to AI inference workloads, yielding up to $15\%$ speedup in decodes over NCCL for a popular LLM, Llama2-70b (Touvron et al., 2023).
+2. Background and Motivation
+2.1. Collective Communication for AI
 Modern AI applications are usually based on large machine learning (ML) models that require training and serving on multiple accelerator devices (GPUs in this paper) in a collective manner.
@@
 This distribution requires GPUs to share intermediate results at some points during the computation, which is done using collective communication operations.
-AllReduce is a common collective communication operation that sums up the partial results from all GPUs and broadcasts the computed result to all GPUs.
+AllReduce is a common collective communication operation that sums the partial results of all GPUs and broadcasts the computed result to all GPUs.
 AllReduce can be divided into two other collective operations: ReduceScatter and AllGather.
-ReduceScatter sums input buffers on all GPUs and distributes the output buffer equally on all GPUs.
+ReduceScatter sums the input buffers on all GPUs and distributes the output buffer equally on all GPUs.
 AllGather collects a distributed buffer from all GPUs and stores the full buffer on each GPU.
 ReduceScatter and AllGather are often used separately depending on the ML model architectures.
-2.2 Limitations of Existing Abstractions
-Existing collective communication libraries internally implement conventional networking abstractions such as send and recv inside GPU kernels.
-This would be a straightforward approach to generalize arbitrary communication algorithms in a hardware-agnostic manner.
-However, this would not be the best choice for performance, because such software abstractions are not well aligned with the GPU hardware abstractions.
-Instead, this paper proposes an API design from the perspective of GPU programming, which makes it more straightforward to efficiently utilize the GPU hardware.
-To be specific, we provide an overview of the NCCL architecture in Section 2.2.1 and discuss its inefficiencies in Section 2.2.2.
-This would be a representative example of existing collective communication libraries for GPUs, because other popular libraries, such as ROCm Collective Communication Library (RCCL) [4] and Microsoft Collective Communication Library (MSCCL) [9], are designed based on NCCL and share the same limitations.
-2.2.1 NCCL Architecture
+2.2. Existing Hardware Abstractions
+Existing collective communication libraries implement conventional networking abstractions such as send and recv inside GPU kernels.
+This is a straightforward approach to generalize arbitrary communication algorithms in a hardware-agnostic manner.
+However, it is not the best choice for performance because such software abstractions are not well aligned with the GPU hardware abstractions.
+Instead, this paper proposes an API design from the perspective of GPU programming, which makes it easier to efficiently utilize GPU hardware.
+In this section, we provide an overview of the NCCL architecture and discuss its inefficiencies.
+This is representative of existing collective communication libraries for GPUs, because other popular libraries, such as RCCL (AMD, 2025) and MSCCL (Cowan et al., 2022), are designed based on NCCL and share the same limitations.
+2.2.1. NCCL Architecture
 NCCL provides a C++ API to setup connections between multiple GPUs and to run communication operations over GPUs.
 All participating processes (that usually hold one GPU each) of a distributed ML application initialize a NCCL handle to build connections beforehand, and then use the handle to call NCCL kernels (which perform collective operations on GPUs) during the execution.
-ML frameworks, such as PyTorch [27] and TensorFlow [2], wrap the NCCL API in Python, which are called by an ML model.
+ML frameworks, such as PyTorch (Paszke et al., 2019) and TensorFlow (Abadi et al., 2016), wrap the NCCL API in Python, which are called by an ML model.
 1global ringRS(in, nelem, ring)
@@
 24 prims.reduce(recv, in+idx, sz, "+");
-Figure 1: Ring ReduceScatter kernel in NCCL.
-Figure 2: Workflow of the NCCL send and recv primitives between two GPUs connected using InfiniBand (IB).
-Even though IB NICs can directly read and write GPU memory, we need the CPU to initiate the transfer (ibv_post_send) and wait for the completion (ibv_poll_cq).
-GPU 0 waits before calling send for next batch of data, and GPU 1 waits before the receive buffer is ready.
+Figure 1. Ring ReduceScatter (simplified) in NCCL.
 NCCL Initialization.
-During initialization, NCCL obtains following information: (i) the rank of each process, (ii) number of processes, (iii) number of distributed nodes in the cluster, (iv) number of GPUs per node, (v) all links between GPUs in a node, including PCIe, NVLink, and xGMI, and (vi) all links between nodes, like InfiniBand (IB) and Ethernet.
+During initialization, NCCL obtains following information: (i) the rank of each process, (ii) number of processes, (iii) number of distributed nodes in the cluster, (iv) number of GPUs per node, (v) all links between GPUs in a node, including PCIe, NVLink, and xGMI, and (vi) all links between nodes, like IB and Ethernet.
 NCCL uses this information to create communication topologies such as ring and tree with bi-directional links.
 Then, NCCL allocates a send buffer that stores data to send to destination GPU and a receive buffer that stores data received from a source GPU.
+Instead, users can also register their own send and receive buffers (using ncclMemAlloc()).
 The metadata of the send and receive buffers on a GPU are shared with all peer GPUs, so that GPUs can either peer-to-peer access buffers of another GPU (through PCIe, NVLink, or xGMI) or use external hardware to copy data (through IB or Ethernet).
@@
 NCCL kernels are built using four building block operations, called primitives, which are send, recv, copy, and reduce.
-(i) send copies data from the send buffer of source GPU to the receive buffer of destination GPU, (ii) recv waits until the transfer by source GPU’s send primitive has finished execution and unblocks send of the source GPU, (iii) copy copies source to destination buffer on the same GPU using GPU threads, and (iv) reduce does elementwise operation, like addition, on two buffers on the same GPU and write the result to another buffer.
-Subsequent calls to send blocks until destination GPUs have called recv to ensure that it is safe to write to send buffer.
-Additionally, NCCL defines extra operations that performs multiple of primitives in a single function call but that does not improve over the limitations we discuss in Section 2.2.2.
+(i) send copies data from the send buffer of source GPU to the receive buffer of destination GPU, (ii) recv waits until the transfer by source GPU’s send has finished execution and unblocks send of the source GPU, (iii) copy copies source to destination buffer on the same GPU using GPU threads, and (iv) reduce does elementwise operation, like addition, on two buffers on the same GPU and write the result to another buffer.
+Calls to send blocks until destination GPUs have called recv to ensure that it is safe to write to send buffer.
+Additionally, NCCL defines extra operations that execute multiple primitives in a single fused function.
+However, this does not improve over the limitations we discuss in Section 2.2.2.
 As an example, Figure 1 describes the NCCL kernel of a ring-based ReduceScatter for half precision floating point (FP16) using the NCCL primitives.
-[^4] The kernel has been simplified for readability.
 The kernel takes the input buffer, number of elements, a ring of GPUs, and writes output to the input buffer.
@@
 The main loop of the kernel performs ReduceScatter in a batch of send/receive buffer sizes (lines 724).
-The loop first copies each batch of input data and send it to the next GPU (lines 911).
-Then the loop goes through all other ranks in the ring and for each rank (i) computes the corresponding offset in the input buffer for the rank, (ii) receives the partially reduced data from the previous GPU in the receive buffer, (iii) reduces this received data with the input offset and stores in the send buffer, and (iv) sends this data to the next GPU in the rank (lines 1419).
-Finally, the GPU stores the data for offset on this GPU (line 24).
-2.2.2 Limitations of NCCL Primitives
+The loop first copies each batch of input data and sends it to the next GPU (lines 911).
+The loop then goes through all other ranks in the ring and for each rank (i) computes the corresponding offset in the input buffer for the rank, (ii) receives the partially reduced data from the previous GPU in the receive buffer, (iii) reduces this received data with the input offset and stores it in the send buffer, and (iv) sends this data to the next GPU in the rank (lines 1419).
+Finally, the GPU stores the data for the offset on this GPU (line 24).
+2.2.2. Limitations of NCCL Primitives
 The send and recv primitives are not flexible enough to utilize a single-instruction multiple-thread (SIMT) architecture like GPUs efficiently.
@@
 Otherwise, it may underutilize the highly-parallel cores and the large memory bandwidth.
-GPU kernels usually case-by-case optimize the workload to carefully distribute tasks among threads and minimize synchronization overhead.
+GPU kernels usually optimize the workload case-by-case to carefully distribute tasks among threads and minimize synchronization overhead.
 Unfortunately, NCCL only implements a static abstraction that groups 128 - 640 threads to collectively call a single primitive and synchronize them at the end of the call.
@@
 An alternative would be making the send and recv asynchronous, which is also against the existing abstractions in NCCL.
+Inflexible Synchronization.
+To ensure data consistency, the NCCL primitives are designed to be self-synchronous.
+Such synchronizations are often too conservative to realize correct communication semantics.
+Consider a loop where $N$ GPUs produce their own data, and each GPU consumes data from all other GPUs, and starts all over again.
+When a single buffer is written by producers and read by consumers, two types of cross-GPU barriers are required: after each producer to ensure the following remote consumer sees consistent data, and after each consumer to prevent the following remote producer from overwriting what is being read by this consumer.
+To reduce synchronization overhead, we can use two buffers in rotation and only keep the first type of barriers which is sufficient to prevent overwriting.
+However, this is not possible with self-synchronized NCCL primitives.
+Figure 2.
+Workflow of the NCCL send and recv primitives between two GPUs connected using InfiniBand (IB).
+Even though IB NICs can directly read and write GPU memory, we need the CPU to initiate the transfer (ibv_post_send) and wait for the completion (ibv_poll_cq).
+GPU 0 waits before calling send for next batch of data, and GPU 1 waits before the receive buffer is ready.
 Interconnect Optimizations.
@@
 For example, intra-node links, like PCIe and NVLink, have two modes of data transfer: (i) thread-copy, where multiple GPU threads read from source GPU memory and write to destination GPU memory, and (ii) DMA-copy, where the CPU initiates the DMA engine of the GPU to copy from source memory to destination memory using cudaMemcpyDeviceToDevice.
-Comparing to that of DMA-copy, thread-copy achieves lower latency, but often utilizes less bandwidth.
+Compared to DMA-copy, thread-copy achieves lower latency, but often utilizes less bandwidth.
 For example, in our AllGather experiments over 8 NVIDIA A100 80G GPUs, thread-copy achieves only up to 227 GB/s NVLink bandwidth, while DMA-copy achieves 263 GB/s (+15.8%).
@@
 [^4] NCCL uses DMA copy only when peer-to-peer access between GPUs is not supported by the hardware configuration.
-Moreover, using DMA-copy frees GPU threads to other work, such as computations, which can lead to better overlapping of computation and communication depending on the application.
-Inflexible Synchronization.
-NCCL primitives are self-synchronized to ensure data consistency.
-However, such synchronizations are often too conservative to realize correct communication semantics.
-Consider a loop where $N$ GPUs produce their own data, and each GPU consumes data from all other GPUs, and start all over again.
-When a single buffer is written by producers and read by consumers, two types of cross-GPU barriers are required: after each producer to ensure the following remote consumer sees consistent data, and after each consumer to prevent the following remote producer from overwriting what is being read by this consumer.
-To reduce synchronization overhead, we can use two buffers in rotation and only keep first the type of barriers which still ensures consistency and prevents overwriting, but is not possible with self-synchronized NCCL primitives.
-Co-optimization with Computation.
-The most common practices in ML frameworks call separate GPU kernels for computation (e.g., cuBLAS or cuDNN) and collective communication (e.g., NCCL).
-However, many recent works [29, 15, 12, 17, 8, 39, 37] have shown performance gains by breaking the barrier and co-optimizing computation and communication operations.
-Unfortunately, coupling NCCL kernels with computation kernels is challenging because NCCL does not design the send and recv primitives as a programmable interface.
-Since the primitives are blocking and use a FIFO queueing mechanism, making them directly work with other parts of the application is not straightforward and necessitates a memory barrier in between.
-A lower-level GPU communication abstraction would be needed to enable the co-optimizations.
-2.3 Limitations of Existing User Interfaces
+Moreover, using DMA-copy frees GPU threads to do other work.
+2.3. Existing Communication Libraries
 NCCL API is designed to provide end-to-end implementations of collective communication operations.
 This would be what most applications expect from the library, but this is fundamentally limited to support diverse performance-sensitive applications.
-Like other GPU workloads do, collective communication over GPUs should be differently optimized for different scenarios, depending on the input size, GPU architecture, networking topology, and the application characteristics [32, 12].
+Like other GPU workloads do, collective communication over GPUs should be differently optimized for different scenarios, depending on the input size, GPU architecture, networking topology, and the application characteristics (Shah et al., 2023; Hwang et al., 2023a).
 It is very difficult for the library developers to provide the best performance for all cases in a timely manner.
-MSCCL [9] is a previous work that tackles this issue by allowing developers to implement their own custom communication algorithms using a DSL.
+MSCCL (Cowan et al., 2022) tackles this issue by allowing developers to implement their own custom communication algorithms using a DSL (Cowan et al., 2023).
 However, MSCCL is still based on the NCCL primitives and does not provide enough flexibility to fully utilize the hardware resources, as described in Section 2.2.2.
-The MSCCL++ Primitive API we propose in this paper is designed at a lower level that barely abstracts the GPU hardware, which enables optimization in the hardware’s perspective.
-This would provide a more flexible and efficient way to optimize GPU communication.
-Inspired by MSCCL, we also provide a DSL for software developers to easily develop communication algorithms.
-3 MSCCL++ Overview
-Figure 3: Overview of MSCCL++.
-3.1 Hierarchical User Interfaces
-MSCCL++ provides three user APIs at hierarchical levels of communication abstractions, as shown in Figure 3: Primitive, DSL, and Collective APIs.
+3. MSCCL++ Overview
+Figure 3. Overview of the MSCCL++ stack.
+3.1. Hierarchical User Interfaces
+MSCCL++ provides three user APIs at hierarchical levels of communication abstractions: Primitive, DSL, and Collective APIs (see Figure 3).
 The three levels provide different trade-offs in terms of programming effort or expertise and execution latency or performance.
 The closer the interface is to the hardware, the higher the programmer’s control on execution, so higher the performance and programming effort.
-MSCCL++ Primitive API (or MSCCL++ API).
-This replaces NCCL’s primitive API with a set of core communication functions that are called from the GPU kernels.
+MSCCL++ Primitive API.
+This replaces NCCL’s primitives with a set of core communication functions that are called from the GPU kernels.
 This is paired with the bootstrapping API called from the host (CPU) side, which is used for connection setup between GPUs.
@@
 Programmers can also use the primitive interface directly to implement application-specific or hardware-specific optimizations in their own GPU kernels, including those kernels that also perform computation.
-As the primitive interface is our key contribution, we focus on its details throughout Section 3 and 4.
+As the primitive interface is our key contribution, we focus on its details in Section 3 and 4.
 MSCCL++ DSL API.
-We reimplement the MSCCLang [10] DSL over MSCCL++, which allows to write custom collective communication algorithms in a high-level language.
-This interface is aimed at users who want to generate communication kernels that are optimized for their own workloads (application and hardware).
-We elaborate in Section 4.3.
+We reimplement the MSCCLang (Cowan et al., 2023) DSL over MSCCL++, which allows us to write custom collective communication algorithms in a high-level language.
+This interface is aimed at users who want to specify communication algorithms or schedules and generate kernels that are optimized for their own workloads (application and hardware).
+The generated kernel uses the Primitive API by running the DSL Executor, so it may have small runtime overheads.
 MSCCL++ Collective API.
@@
 This library has the same limitations as that of NCCL (Section 2.3); i.e., it may not provide the best algorithm for certain workloads.
-For better performance, users can install their own optimized algorithms written using the MSCCL++ DSL API into this library.
-3.2 Primitive Communication Abstractions
-3.2.1 Communication Channels
-Since the mode(s) of data transfer supported differs between interconnects, MSCCL++ defines an abstract channel for each data transfer mode supported by the hardware and exposes communication primitives specific to that channel that can be invoked directly from inside a GPU kernel.
-MSCCL++ defines three communication channels [^7] For readers who are familiar with NCCL: MSCCL++ channel does not correspond to that in NCCL.: PortChannel, MemoryChannel, and SwitchChannel, which correspond to port-mapped I/O, memory-mapped I/O, and switch-mapped I/O respectively.
-A PortChannel uses interconnect ports to communicate between GPUs; i.e., a GPU can initiate the data transfer through a port to another GPU (note that ports are controlled by dedicated hardware for I/O, such as DMA engines on GPUs or RDMA NICs).
+For AI inference workloads, we implement the best algorithms in our collective kernels using the SpeCL Primitive API.
+3.2. Primitive Communication Abstractions
+3.2.1. Communication Channels
+The supported mode(s) of data transfer differs between interconnects.
+Therefore, MSCCL++ defines an abstract channel for each data transfer mode supported by the hardware, and exposes communication primitives specific to that channel that can be invoked directly inside a GPU kernel.
+MSCCL++ defines three communication channels (for readers who are familiar with NCCL, MSCCL++ channel does not correspond to that in NCCL): PortChannel, MemoryChannel, and SwitchChannel, which correspond to port-mapped I/O, memory-mapped I/O, and switch-mapped I/O respectively.
+A PortChannel uses interconnect ports to communicate between GPUs; i.e., a GPU can initiate the data transfer through a port to another GPU.
+Ports are controlled by dedicated hardware for I/O, such as DMA engines on GPUs or RDMA NICs.
 A MemoryChannel uses peer-to-peer memory access to communicate between GPUs; i.e., a GPU can directly access another GPU’s memory.
 A SwitchChannel uses interconnection-switch-enabled multimem memory access to communicate between GPUs; i.e., a GPU can access the memory of multiple GPUs simultaneously through a switch.
-The same interconnect can support multiple modes of data transfer, so multiple channels can be supported on the same interconnect; for example, all three channels are supported by NVLink, only PortChannel and MemoryChannel are supported by xGMI and PCIe, while only PortChannel is supported by InfiniBand.
+The same interconnect can support multiple modes of data transfer, so multiple channels can be supported on the same interconnect; e.g., all three channels are supported by NVLink, only PortChannel and MemoryChannel are supported by xGMI and PCIe, while only PortChannel is supported by InfiniBand.
 The implementation of each channel is specific to the interconnect.
@@
 On the other hand, the PortChannel uses the DMA-copy data transfer mode over NVLink, xGMI, PCIe, and Infiniband.
-Current hardware interconnects require a CPU thread to initiate the data transfer.
-For example, transferring data over InfiniBand requires the CPU to call ibv_post_send that initiates the RDMA transfer from the sending GPU memory to the receiving GPU memory.
-However, such details are handled transparently by the implementation.
+Current interconnects require a CPU thread to initiate the data transfer; e.g., transferring data over InfiniBand requires the CPU to call ibv_post_send which initiates the RDMA transfer from the sending GPU memory to the receiving GPU memory.
+Such details are handled transparently by the implementation.
 Generality.
@@
 For example, if future hardware interconnects supports initiating DMA-copy from within a GPU kernel, the same PortChannel API can target the new hardware.
+3.2.2. Communication Primitives
+All MSCCL++ primitives are defined as a method of a channel.
+The primitives are all conceptual and do not imply any specific implementation.
+In this section, we describe the primitives API for the PortChannel with an example.
+Primitives for the other channels, and the implementation for all the channels are explained in Section 4.
+Borrowing the term from MPI (Snir et al., 1998), writing data to the peer’s side through a channel is called put.
+To provide our communication abstraction as close as possible to hardware capabilities, we design put to be zero-copy (i.e., no intermediate buffers), one-sided (i.e., initiated by a peer without participation of the other peer), and asynchronous.
+MSCCL++ provides other interfaces for synchronization purposes.
+Specifically, signal and wait provide a mechanism for synchronization (including memory consistency) across GPUs.
+flush is for local synchronization that ensures all previous put operations are already on the fly.
+Figure 4 shows the semantics of four MSCCL++ primitives: put, signal, wait, and flush.
+The put primitive transfers data from src0 buffer of GPU-0 to dst1 buffer of GPU-1.
+GPU-0 calls a following signal primitive that signals GPU-1, which is asynchronous yet strictly ordered with the previous put.
+GPU-1 has to call the wait primitive before it can read dst1.
+GPU-0 calls the flush primitive to ensure that the previous put is on the fly and src0 can be safely reused.
+MSCCL++ also supports a few fused primitives that can reduce the overhead of API calls.
+For example, since a put is usually followed by a signal, we provide the putWithSignal function that conducts both at once.
 1//async
@@
 (b) GPU-1
-Figure 4: MSCCL++ data transfer abstractions.
+Figure 4.
+MSCCL++ data transfer abstractions.
 put asynchronously transfers data from one GPU to another.
@@
 flush ensures the completion of preceded data transfer.
-3.2.2 Communication Primitives
-All MSCCL++ primitives are defined as a method of a channel.
-The primitives are all conceptual and do not imply any specific implementation.
-In this section, we describe the primitives API for the PortChannel with an example.
-Primitives for the other channels, and the implementation for all the channels are explained in Section 4.
-Borrowing the term from MPI [34], writing data to the peer’s side through a channel is called put.
-To provide our communication abstraction as close as possible to hardware capabilities, we design put to be zero-copy (i.e., no intermediate buffers), one-sided (i.e., initiated by a peer without participation of the other peer), and asynchronous.
-MSCCL++ provides other interfaces for synchronization purposes.
-Specifically, signal and wait provide a mechanism for synchronization (including memory consistency) across GPUs.
-flush is for local synchronization that ensures all previous put operations are already on the fly.
-Figure 4 shows the semantics of four MSCCL++ primitives: put, signal, wait, and flush.
-The put primitive transfers data from src0 buffer of GPU-0 to dst1 buffer of GPU-1.
-GPU-0 calls a following signal primitive that signals GPU-1, which is asynchronous yet strictly ordered with the previous put.
-GPU-1 has to call the wait primitive before it can read dst1.
-GPU-0 calls the flush primitive to ensure that the previous put is on the fly and src0 can be safely reused.
-MSCCL++ also supports a few fused primitives that can reduce the overhead of API calls.
-For example, since a put is usually followed by a signal, we provide the putWithSignal function that conducts both at once.
+Example.
+Figure 5 shows MSCCL++ implementation of the all-pairs ReduceScatter algorithm, which is not used in NCCL.
+In this algorithm, all GPUs send their data at $\frac{i}{N}$ offset to GPU $i$, the $i^{th}$ GPU does elementwise reduction of the received data, and stores the reduced data.
+To store the received data, the implementation allocates a scratch buffer per GPU.
+The kernel’s main loop performs reduction in batches of the scratch size (line 518).
+On each GPU, the loop first put $\frac{1}{N}$ of data to all other GPUs’ scratch buffer and then signals the completion of transfer (line 910).
+Then a GPU waits for the data to arrive from other GPUs and reduce it with its part of input data (line 1315).
+Finally, a barrier among all GPUs is required to ensure scratch is not overwritten for next main loop offset (line 18).
+Section 5 shows that this algorithm works better than the ring algorithm for smaller data sizes while the ring algorithm works better for large data sizes.
+3.2.3. Advantages of Communication Primitives
 1global allPairsRS(count, gpus, channels[gpus])
@@
 11
-12 //Reduce ever pair of GPU
+12 //Reduce every pair of GPU
 13 for (g = 0; g < gpus.num-1; g++)
@@
 18 multiDeviceBarrier();
-Figure 5: All-pairs ReduceScatter kernel in MSCCL++. Channels are initialized with source as input and destination as scratch buffer.
-Example.
-Figure 5 shows MSCCL++ implementation of the all-pairs ReduceScatter algorithm, which is not used in NCCL.
-In this algorithm, all GPUs send their data at $\frac{i}{N}$ offset to GPU $i$, the $i^{th}$ GPU does elementwise reduction of the received data, and stores the reduced data.
-To store the received data, the implementation allocates a scratch buffer per GPU.
-The kernel’s main loop performs reduction in batches of the scratch size (line 518).
-On each GPU, the loop first put $\frac{1}{N}$ of data to all other GPUs’ scratch buffer and then signals the completion of transfer (line 910).
-Then a GPU waits for the data to arrive from other GPUs and reduce it with its part of input data (line 1315).
-Finally, a barrier among all GPUs is required to ensure scratch is not overwritten for next main loop offset (line 1818).
-Section 5 shows that this algorithm works better than the ring algorithm for smaller data sizes while the ring algorithm works better for large data sizes.
-3.3 Advantages
+Figure 5. All-pairs ReduceScatter kernel in MSCCL++. Channels are initialized with source as input and destination as scratch buffer.
 The MSCCL++ Primitive API itself does not introduce any novel optimization techniques – instead, it offers new straightforward abstractions that enable quick optimization of GPU communication.
-The key advantage comes from exposing the primitive functionalities as a user interface, which introduces three benefits.
-First, it allows users to easily customize GPU communication algorithms for their own infrastructure and workloads (see following paragraphs).
-This is valuable in practice because the library developers may not be aware of the specific workloads and hardware configurations of their users, hence the difficulties of timely support of efficient implementations.
-Second, the library developers can quickly deliver new hardware support by leaving the algorithm design to the users who can do it better (see Section 4.5.1).
-Third, since the primitive interface is close to the hardware, it is more flexible to adopt new hardware features.
-We carefully optimized GPU communication in the MSCCL++ DSL Executor and Collective kernels by leveraging the features in the MSCCL++ Primitive API.
+The key advantages come from exposing the primitive functionalities as a user interface.
+First, it allows application developers to easily and quickly optimize GPU communication algorithms for their own infrastructure and workloads.
+This is valuable in practice because the library developers may not be aware of the specific workloads and hardware configurations of their users.
+Second, the library developers can quickly deliver new hardware support by implementing the same primitive API for the new hardware.
+Third, since the primitive API is close to the hardware, it is more flexible to extend or adopt it for new hardware features.
+MSCCL++ carefully optimizes GPU communication in the DSL Executor and collective kernels leveraging some of the below features in the Primitive API.
 Users of the Primitive API can also implement these optimizations directly.
-Section 5 shows how these optimizations help MSCCL++ obtains better performance than NCCL.
 Asynchronous Communication.
-MSCCL++ contains separate primitives for data transfer and synchronization, thus, enabling asynchronous communication and batching the synchronization of multiple communication into a single synchronization.
-This abstraction has three benefits: (i) allows more algorithms to be specified than the synchronous communication in NCCL, (ii) enables faster collective communication performance by allowing an algorithm space that can batch several synchronizations into one, (iii) frees up resources of peer GPUs to do other meaningful work inside the application, and (iv) allows better inter-kernel optimizations.
+MSCCL++ contains separate primitives for data transfer and synchronization, allowing asynchronous communication and batching synchronization of multiple data transfers.
+This abstraction has multiple benefits: (i) allows more algorithms to be specified than the synchronous communication in NCCL, (ii) enables faster collective communication performance by enabling algorithms to combine several synchronizations into one, (iii) frees up resources of peer GPUs for other meaningful work of the application, and (iv) enables better inter-kernel optimizations.
 Specialized GPU Kernels.
@@
 In contrast, by using different channel types in MSCCL++ (PortChannel or MemoryChannel), user can utilize both thread-copy and DMA-copy data transfer mode over NVLink.
-Computation-Communication Optimizations.
-Since MSCCL++ Primitive API is an in-kernel interface for communication, we can co-optimize both computation and the communication kernels, leading to better application performance.
-One such optimization is fusing communication with a GeMM kernel [15, 37, 17, 8, 39], so that the number of memory reads/writes decreases, and a GPU’s warp context switching enables overlapping of communication with the GeMM kernel.
-MSCCL++ Primitive API is an easy-to-use and efficient interface to implement this idea, and we remain the implementation as a future work.
-4 Implementation
-4.1 Initialization
-In general, an MSCCL++ program runs as a multi-process on all nodes and GPUs within a node, such that each process is responsible for one GPU.
-Since MSCCL++ enables application-specific optimization, it provides CPU side interfaces for custom initialization of the collective communications used in an application.
-MSCCL++ provides a default initialization process that can be specialized by users for their application.
-The initialization process involves setting up connections between GPUs, finding the topology of GPU connections, and constructing channels and sharing scratch buffers between peer GPUs.
-The first step in initialization is to create a communication channel between all distributed CPU processes to exchange metadata between processes.
-We call this communication channel a Bootstrap.
-A bootstrap consists of four virtual methods: (i) send to send data from one CPU process to another, (ii) recv to recv data from a CPU process, (iii) allGather to perform the AllGather collective communication, and (iv) barrier to synchronize all distributed processes.
-MSCCL++ implements the default bootstrap methods using POSIX sockets.
-If a user prefers another distributed communication protocol such as MPI or torch.distributed, it is straightforward to override bootstrap methods because all these methods directly correspond to existing distributed protocols.
-The second step for application is to construct a communicator object using the bootstrap object.
-The application uses the communicator object to register buffers that are used to perform data transfers.
-These buffers can be either application buffers themselves or scratch buffers dedicated for communications.
-The communicator then creates relevant channel objects (PortChannel, MemoryChannel, or SwitchChannel) between GPUs based on physical links.
-For each GPU, these channel objects are initialized with buffers.
-Finally, these channels are shared among processes of each GPU.
-4.2 Primitives for Channels and Protocols
-A connection between GPUs in MSCCL++ is called a channel.
-PortChannel and MemoryChannel are peer-to-peer connections (i.e., between 2 GPUs), whereas SwitchChannel is a connection among a group/collection of (2 or more) GPUs.
-Figure 6 shows three types of channels and its interface in MSCCL++.
-A channel is initialized with the source and destination buffers, and an integer allocated on the GPU memory that servers as semaphore.
-In this section, we describe all three kinds of MSCCL++ channels.
 1class Channel
@@
 33 ulong sz, uint tid, uint tids)
-Figure 6: PortChannel, MemoryChannel, and SwitchChannel primitives in MSCCL++. Implementation of MemoryChannel primitives differ for the type of protocol: LL or HB.
-4.2.1 PortChannel
-A PortChannel implements primitives when data transfer is done over ports connected to GPU memory, such as using InfiniBand API’s ibv_post_send for IB or cudaMemcpy for DMA copy.
-Since data transfer over a port require the CPU to initiate the transfer, this channel creates a CPU thread for each GPU.
+Figure 6. PortChannel, MemoryChannel, and SwitchChannel primitives in MSCCL++. Implementation of MemoryChannel primitives differ for the type of protocol: LL or HB.
+4. SpeCL Design and Implementation
+4.1. Initialization of SpeCL
+In general, a MSCCL++ program runs as a multi-process on all nodes and GPUs within each node, so that each process runs one GPU.
+MSCCL++ provides host (CPU) side interfaces for initializing multiple processes, and the interfaces can be specialized by users for their application.
+The initialization involves setting up connections between GPUs, finding the topology of GPU connections, and constructing channels and buffers.
+The first step in initialization is to exchange metadata between host processes, which we call a Bootstrap.
+It consists of four virtual methods: (i) send to send metadata to another process, (ii) recv to receive metadata from a process, (iii) allGather to collect metadata of all processes, and (iv) barrier to synchronize all processes.
+MSCCL++ implements a default bootstrap methods using POSIX sockets, while it is straightforward to override with other popular implementations such as MPI or torch.distributed.
+The second step is to construct a Communicator that registers communication buffers (either user buffers or dedicated ones) and creates corresponding MSCCL++ channels between GPUs based on the underlying physical links.
+4.2. MSCCL++ Primitive API and Channels
+A MSCCL++ channel is a connection between two or more GPUs.
+PortChannel and MemoryChannel are peer-to-peer (i.e., between two GPUs), whereas SwitchChannel is for a group of (i.e., two or more) GPUs.
+Figure 6 shows three types of channels and its interface in MSCCL++.
+A channel is initialized with the source and destination buffers, and an integer allocated on the GPU memory that servers as a semaphore.
+4.2.1. PortChannel
+A PortChannel implements primitives when data transfer is done over ports connected to GPU memory, such as cudaMemcpy for intra-node DMA or ibv_post_send for RDMA.
+Since data transfer over a port currently requires the CPU to initiate the transfer, each channel creates its own CPU thread.
 The CPU thread reads data transfer or synchronization requests from a first-in-first-out request queue.
-The storage, head, and tail of a request queue are allocated using cudaMallocManaged, so that both CPU and GPU can access this data.
-By default we allocate a queue of 1024 requests but this size can be changed by the user.
+The storage, head, and tail of a request queue are allocated using cudaMallocManaged as both CPU and GPU can access it.
 We now discuss the workflow of PortChannels using Figure 7.
-Figure 7: PortChannel workflow for IB starting from ⓪ when GPU 0 calls put primitive to ⑧ when GPU 1 receives the data.
-0
-When the GPU calls a primitive, such as, put, then the first participating thread of the GPU pushes the request to the queue by writing at the head.
+0 When the GPU calls a primitive, such as put, the first participating thread of the GPU pushes the request to the queue by writing at the head.
 Before writing to the queue, the GPU checks if the queue is filled, i.e., if the head value is more than the tail value.
@@
 1 Then the GPU increments the head to the next element.
-2 The CPU thread continously reads the element at the tail to see if there is a request from the GPU and when there is a request, then the CPU reads the request, zeros out the current element, and increments the tail.
+2 The CPU thread continuously reads the element at the tail to see if there is a request from the GPU, and when there is a request, the CPU reads the request, zeros out the current element, and increments the tail.
 3 Now the CPU thread will process the request.
 Below we explain the requests generated by primitives and how the CPU handle these requests.
+Figure 7. PortChannel workflow for IB from ⓪ when GPU 0 calls put primitive to ⑧ when GPU 1 receives the data.
 Data Transfer.
@@
 Flush.
-Primitive flush pushes a flush request in the queue and the first thread of the GPU wait until the queue head is equal to or more than the queue tail.
-The CPU process the flush request by waiting until all the previous data transfer and synchronization requests have been completed.
+Primitive flush pushes a flush request in the queue and the first thread of the GPU waits until the queue head is equal to or more than the queue tail.
+The CPU processes the flush request by waiting until all the previous data transfer and synchronization requests have been completed.
 For example, for IB we use ibv_poll_cq to get the status of all requests.
 After the flush request is complete, the GPU is free to re-write to the source buffer.
-4.2.2 MemoryChannel
+4.2.2. MemoryChannel
 A MemoryChannel wraps data transfer methods that use thread-copy mechanism, i.e., directly use GPU threads for writing to peer GPUs memory.
@@
 Both protocols achieves these properties by synchronizing on the suitable granularity of data.
-(i) In HB protocol, peer GPUs transfer a large chunk of data and synchronize this chunk once using wait and signal primitives.
+(i) In the HB protocol, peer GPUs transfer a large chunk of data and synchronize this chunk once using wait and signal primitives.
 Hence, the synchronization time is amortized over the transfer time leading to high bandwidth.
-However, since the receiving GPU needs to wait for the whole chunk before the GPU can process it, the HB protocol also has high latency.
+However, since the receiving GPU needs to wait for the whole chunk before processing it, the HB protocol also has high latency.
 (ii) In LL protocol, the receiving GPU synchronizes on fewer chunks of data being transferred and process the chunk as soon as it is received, thus, achieving lower latency.
@@
 The put primitive requires an extra integer flag value.
-For every $N-1$ elements written to the receiving GPU, put also write the flag.
-The receiving GPU uses the read primitive waits until the flag value at $N$ index of receiving buffer is set and then read and return the $N-1$ elements.
+For every $N-1$ elements written to the receiving GPU, put also writes the flag.
+The receiving GPU uses the read primitive, waits until the flag value at $N$ index of receiving buffer is set, and then reads and returns the $N-1$ elements.
 However, we cannot arbitrarily use any $N$ because GPUs follow a weak memory consistency model where writes to different memory locations by multiple threads can be performed in any order.
@@
 A protocol defines the implementation of data transfer and synchronization primitives.
-We now describe the implementation of primitives based on protocol as follows:
+We now describe the implementation of primitives based on the protocol as follows:
 Data Transfer.
 Primitive put reads data elements from source buffer and writes to destination buffer starting from different offsets.
-Similarly, the read primitive reads and return data from the destination buffer, and the write primitive writes data including the flag for the LL protocol to the destination buffer.
+Similarly, the read primitive reads and returns data from the destination buffer, and the write primitive writes data including the flag for the LL protocol to the destination buffer.
 To maximize bandwidth for the HB protocol, both primitives use 16-bytes loads and stores.
 For the LL protocol, both primitives also takes a flag and uses 8-byte load and stores by default or user supplied vector length.
-Since put both primitives are called by multiple threads (all or first few threads of the kernel), we achieve maximum bandwidth.
+Since both put primitives are called by multiple threads (all or first few threads of the kernel), we achieve maximum bandwidth.
 Synchronization.
-Synchronization primitives wait on a semaphore which is an integer allocated on the receiving GPU.
-The signal primitive atomically increments the semaphore of the receiving GPU and calls threadfence_system to ensure that writes by put and semaphore increments are made available in this order.
+Synchronization primitives wait on a semaphore that is an integer allocated on the receiving GPU.
+The signal primitive atomically increments the receiving GPU’s semaphore and calls threadfence_system to ensure that writes by put and semaphore increments are made available in this order.
 The wait primitive performs a busy-wait while-loop that checks until the value of semaphore has reached the expected value.
@@
 The channel tracks the expected value using its expectedValue member.
-Since after put returns memory writes are already in operation, src buffer can be reused.
-Therefore, the flush primitive is empty.
-4.2.3 SwitchChannel
+The flush primitive is a no-op because after put returns, the source buffer can be reused, even though the write is still in progress.
+4.2.3. SwitchChannel
 A SwitchChannel provides primitives for performing collective operations among GPUs.
@@
 A SwitchChannel in MSCCL++ provides two primitives: reduce to add the corresponding elements of buffers residing on different GPUs and broadcast to send elements from a buffer on a GPU to all other GPUs.
-We now discuss the implementation of these primitives for NVIDIA NVSwitch, which provides in-network, switch-based aggregation and multicast capabilities using the NVLink SHARP (NVLS) technology [23].
+We now discuss the implementation of these primitives for NVSwitch, which provides in-network, switch-based aggregation and multicast capabilities using NVLink SHARP (NVLS) technology (NVIDIA, 2025c).
 However, we believe that these implementations can be generalized to other hardware in the future.
@@
 The multimem instruction sends the register value to the switch, which broadcasts and stores the value to all the virtual memory addresses pointed by the multimem address.
-4.3 DSL Implementation
+4.3. MSCCL++ DSL API and Executor
 MSCCL++ DSL API is a Python-based language that allows describing a communication algorithm at a high level.
 It converts the algorithm description into a sequence of instructions that can be executed by the DSL Executor, which is a GPU kernel that reads and runs instructions back-to-back.
-MSCCL++ DSL is an extension of the MSCCLang [10] that supports new instructions based on the MSCCL++ Primitive API and lifts a few restrictions of the language; e.g., we enable a single GPU thread block to access multiple GPUs at the same time, etc.
-MSCCL++ DSL introduces a few limitations over the Primitive API.
-First, as the DSL introduces higher level instructions than the Primitive API, it may not be the most efficient way to implement.
-In our evaluation, DSL-written algorithms perform 3% worse than the hand-written ones on average, and is up to 18% worse in a corner case.
-Second, since a DSL-written algorithm can be run only by the provided executor kernel, it may not be straightforward to merge with other computation kernels if needed.
-Despite limitations, MSCCL++ DSL is still useful for quick prototyping of collective communication and easy-to-understand algorithm description.
-4.4 Collective API and Algorithms
-MSCCL++ Collective API is the highest level of API that MSCCL++ supports, and it is the same as the original NCCL API.
-It is implemented as a wrapper of the MSCCL++ DSL API, which means that users can install a DSL-written algorithm to be called by the NCCL API.
-Users can also use the default algorithms already provided by the API.
-At runtime, for each communication, MSCCL++ automatically selects the best performing algorithm.
-To be specific, we provide high-level description of the AllReduce algorithms we implement using the MSCCL++ DSL API below.
-Each algorithm can be implemented in multiple ways for optimization depending on hardware and message size.
-For example, we implement eight different versions of the two-phase all-pairs (2PA) algorithm explained in a following paragraph, and we omit the details for brevity.
+MSCCL++ DSL is an extension of the MSCCLang (Cowan et al., 2023) that supports new instructions based on the MSCCL++ Primitive API and lifts a few restrictions of the language; e.g., we enable a single GPU thread block to access multiple GPUs at the same time.
+4.4. MSCCL++ Collective API and Kernels
+MSCCL++ Collective API is the simplest API that MSCCL++ offers, and it is the same as the popular NCCL API.
+We implement this API using kernels written in the MSCCL++ Primitive API.
+Users can either use these default kernels or plug-in their own algorithms written using the MSCCL++ DSL or Primitive APIs.
+We now briefly describe a few AllReduce algorithms included in the default kernels to: (i) provide ideas on low-level considerations for implementing high-performance GPU communication kernels, and (ii) to show how low-level interfaces can be useful for customization on different workloads.
+Env. Name GPU Intra-node Link Network
+A100-40G NVIDIA A100 (40G) (8x/node) NVLink 3.0 Mellanox HDR InfiniBand (200 Gb/s, 1x NIC/GPU)
+A100-80G NVIDIA A100 (80G) (8x/node) NVLink 3.0 Mellanox HDR InfiniBand (200 Gb/s, 1x NIC/GPU)
+H100 NVIDIA H100 (8x/node) NVLink 4.0 Quantum-2 CX7 InfiniBand (400 Gb/s, 1x NIC/GPU)
+MI300x AMD MI300x (8x/node) Infinity Fabric Gen 4 Quantum-2 CX7 InfiniBand (400 Gb/s, 1x NIC/GPU)
+Table 1. List of environments used for evaluation.
 1.
@@
 As we use the 1PA algorithm only for very small message collectives within a single node, we implement only a single version of 1PA that uses the LL protocol with MemoryChannel.
-MSCCL++ can implement it much more efficient than baselines by removing unnecessary memory copies and synchronization as described in Section 2.2.2.
+MSCCL++ can implement it much more efficiently than other libraries by removing synchronizations.
 2.
 Two-phase All-pairs (2PA).
-By two-phase, the algorithm splits the AllReduce into two phases: the first for ReduceScatter (each of $N$ GPUs collect and reduce $1/N$ of the data) and the second for AllGather (each GPU broadcasts the reduced data to all other GPUs) [31].
-Two-phase algorithms are more bandwidth-efficient and conducts less reduction than one-phase algorithms.
+By two-phase, the algorithm splits the AllReduce into two phases: the first for ReduceScatter (each of $N$ GPUs collects and reduces $1/N$ of the data) and the second for AllGather (each GPU broadcasts the reduced data to all other GPUs) (Rabenseifner, 2004).
+Two-phase algorithms are more bandwidth-efficient and conduct less reduction than one-phase algorithms.
 In 2PA, the ReduceScatter and AllGather phases are done in the all-pairs manner each.
-We use 2PA for single-node collectives and implement multiple versions with combinations of options: MemoryChannel or PortChannel, LL or HB protocol, and put-based or get-based.
-Using MSCCL++, we can optimize 2PA in various ways that are not possible with baselines.
-For example, for up to a few MBs of messages, we exploit rotating buffers to reduce synchronization at the cost of using more memory space, as explained in Section 2.2.2.
+We use 2PA for single-node collectives and implement multiple variants using PortChannel, MemoryChannel with LL or HB protocol, or SwitchChannel.
+Using MSCCL++, we can optimize 2PA in various ways that are not possible in existing libraries.
+For example, for up to a few MBs of messages, we exploit rotating buffers to reduce synchronization at the cost of using more memory space.
 As another example, we can let a single thread group read data from multiple other GPUs at the same time.
-This allows efficient data reduction comparing to baselines that read data from different GPUs one-by-one, which syncrhonizes each reduction step.
+This allows efficient data reduction compared with other libraries that read data from different GPUs one-by-one, which synchronizes each reduction step.
 3.
@@
 2PH is a two-phase algorithm that performs ReduceScatter and AllGather in a hierarchical manner each.
-We use it for multi-node collectives and implement two versions.
-The first version is for small messages using LL protocol.
+We use it for multi-node collectives and implement two variants.
+The first version is for small messages using the LL protocol.
 Each node conducts local ReduceScatter that splits the data into the number of GPUs in a node.
-This requires to send more data across nodes and introduce redundant reduction, but it is faster for small messages by reducing synchronization steps.
+This requires sending more data across nodes and introduces redundant reduction, but it is faster for small messages by reducing synchronization steps.
 Cross-node communication is done in all-pairs manner.
@@
 The second version of 2PH is for large messages using HB protocol.
-Unlike the first version, to utilize the link bandwidth efficiently, the number of data chunks is the same as the number of GPUs.
-Similar to the first version, it performs all-pairs cross-node communication and local ReduceScatter in a pipelined manner.
-4.
-Two-Phase All-Pairs Multimem (2PAM).
-2PAM is a specialized algorithm for NVIDIA H100 GPUs that leverages the in-network, switch-based aggregation and multicast capabilities by the NVLink SHARP (NVLS) technology [23].
-This algorithm simply walks through all data elements one-by-one in a loop to call two NVLS multimem instructions, one for in-network cumulating one element from all GPUs, and another for broadcasting the cumulated element.
-4.5 Cross-hardware Implementation
-This section introduces our experience in developing MSCCL++ across different hardware platforms and how we reduce development efforts for future hardware.
-4.5.1 Quick Support for New GPUs
-Extending MSCCL++ to support new hardware features is much easier compared to doing so in existing communication libraries, because the primitive interface is directly exposed to the user.
-This design accelerates the overall development because (1) the library developers only need to add a shallow layer of abstraction over low-level code that features key functionalities, and (2) the algorithm developers can write clever optimizations and fine-tuning with precise control of the hardware efficiency.
-We provide an example of supporting the NVIDIA H100 GPU in MSCCL++, which delivers the hardware-based collective operations using NVSwitch (explained in Section 4.2.3).
-Since this feature introduced a very different concept from previous communication interfaces (i.e., PortChannel or MemoryChannel), we developed the SwitchChannel interface to support this feature.
-The development took only 8 weeks for two developers, including learning the basic usage of this feature, abstracting the feature as a new type of channel (i.e., SwitchChannel), and finally developing a new AllReduce algorithm using SwitchChannel that outperforms NCCL/MSCCL by more than 2.2$\times$ in average for message sizes of 1KB - 1GB.
-As another example, we have developed MSCCL++ to support the AMD MI300x GPU, while the previous version supported only NVIDIA GPUs.
-The development took only 7 weeks for one developer: 3 weeks for basic AMD GPU support and 4 weeks to develop new AllReduce algorithms that outperform RCCL/MSCCL for message sizes of 1KB - 1GB.
+This variant also performs all-pairs cross-node communication and local ReduceScatter in a pipelined manner.
+Unlike the first variant, to utilize the link bandwidth efficiently, the number of data chunks is the same as the number of GPUs.
+4.5. Portability and Implementation Effort
+MSCCL++ is designed to be portable.
+Code using MSCCL++ collectives, algorithms using the MSCCL++ DSL, and kernels using the MSCCL++ primitives are portable as long as the underlying MSCCL++ Primitive API supports that hardware.
+Our design makes it easy to add support for new hardware.
+As an example, MSCCL++ initially supported only NVIDIA GPUs; to add support for AMD MI300x GPUs, our implementation took only 7 weeks for one developer: 3 weeks for basic AMD GPU support and 4 weeks to develop new AllReduce algorithms that outperform RCCL/MSCCL for message sizes of 1KB - 1GB.
 It is especially remarkable that the AMD-specific code in MSCCL++ (excluding makefiles and algorithms) is only less than 10 lines of code, while RCCL, the hard-forked NCCL for AMD GPUs, is substantially diverged from NCCL.
-This is possible because the low-level API of AMD GPUs (i.e., HIP) is almost the same as that of NVIDIA GPUs (i.e., CUDA), and the MSCCL++ Primitive API is only a shallow abstraction on top of the low-level API.
+MSCCL++ could achieve this because the low-level API of AMD GPUs (i.e., HIP) is almost the same as that of NVIDIA GPUs (i.e., CUDA), and the MSCCL++ Primitive API is only a shallow abstraction on top of the low-level API.
 This significantly reduces the effort to maintain and develop MSCCL++.
-4.5.2 Managing Memory Consistency
-Besides the proposed abstractions, MSCCL++ also puts particular considerations on its low-level implementation to reduce dependencies on certain hardware, thus reducing technical debt.
-As an example, we introduce our approach toward managing memory consistency.
-Ensuring memory consistency is critical as a communication library.
-In particular, fast-evolving chips and their compilers often consist of various bugs or undocumented behaviors that can lead to breaking memory consistency, which is often hard to detect and fix.
-This is a solid problem that prevents the quick adoption of the latest hardware techniques.
-From our production efforts to tackle this problem, we found that the most effective and efficient way to ensure memory consistency is relying on the atomic operations, rather than leveraging other consistency models provided by hardware or compiler features.
-This makes it easier to blame the hardware or compiler when we find unexpected behaviors, [^8] Indeed, we found a bug of an atomic instruction in CUDA 12.x NVCC, which was reported to and fixed by NVIDIA.
-and also makes MSCCL++ more portable and maintainable.
-For example, we use atomic instructions that comply with the C++11 memory model instead of volatile instructions.
-As another example, for signal over RDMA, one may simply send a separate RDMA write request to send a flag to the target GPU, assuming that the underlying hardware and stacks are configured and working correctly to ensure ordered write.
-However, instead, MSCCL++ uses the atomic fetch-and-add operation of RDMA to ensure consistency from the software perspective (the impact on performance is negligible small).
-Env. Name GPU Intra-node Link Network
-A100-40G NVIDIA A100 (40G) (8x/node) NVLink 3.0 Mellanox HDR InfiniBand (200 Gb/s, 1x NIC/GPU)
-A100-80G NVIDIA A100 (80G) (8x/node) NVLink 3.0 Mellanox HDR InfiniBand (200 Gb/s, 1x NIC/GPU)
-H100 NVIDIA H100 (8x/node) NVLink 4.0 Quantum-2 CX7 InfiniBand (400 Gb/s, 1x NIC/GPU)
-MI300x AMD MI300x (8x/node) Infinity Fabric Gen 4 Quantum-2 CX7 InfiniBand (400 Gb/s, 1x NIC/GPU)
-Table 1: List of environments used for evaluation.
-5 Evaluation
+The MSCCL++ design makes it easy to extend the Primitive API to support new hardware features.
+For example, NVIDIA NVLink 4.0 introduced support for multimem instructions for switch-based aggregation and multicast.
+Since this feature introduced a very different concept from previous interfaces (i.e., PortChannel or MemoryChannel), we developed the SwitchChannel interface to support this feature.
+The development took only 8 weeks for two developers, including learning the basic usage of this feature, abstracting the feature as a new type of channel (i.e., SwitchChannel), and finally developing a new AllReduce algorithm using SwitchChannel that outperforms NCCL/MSCCL by more than $\times2$ in average for message sizes of 1KB - 1GB.
+5. Evaluation
 Environments.
@@
 Each node is equipped with 8 GPUs per node (either A100, H100, or MI300X), intra-node links between GPUs (either NVLink or Infinity Fabric (a.k.a.
-xGMI)), and inter-node InfiniBand links (one NIC per GPU).
-All NICs are connected to a single InfiniBand networking switch.
+xGMI)), and inter-node IB links (one NIC per GPU).
+All NICs are connected to a single IB networking switch.
 NVIDIA GPUs use CUDA 12.4, while AMD GPUs use ROCm 6.2.
@@
 Baselines.
-We compare MSCCL++ with existing collective communication libraries, including NCCL 2.23.4 [21], RCCL 2.20.5 [4], [^9] As RCCL partly adopts MSCCL, to prevent confusion, we refer to MSCCL-disabled RCCL.
-and MSCCL 2.23 [19].
+We compare MSCCL++ with the state-of-the-art collective communication libraries, including NCCL 2.26.2 (NVIDIA, 2025a), RCCL 2.20.5 (AMD, 2025), [^5] As RCCL partly adopts MSCCL, to prevent confusion, we refer to MSCCL-disabled RCCL.
+and MSCCL 2.23 (Microsoft, 2025a).
 MSCCL borrows the stack implementation of NCCL/RCCL and supports custom algorithms.
 Therefore, MSCCL performs exactly the same as NCCL/RCCL if they use the same algorithm.
-5.1 Collective Communication
-We present collective communication performance of MSCCL++ and compare it with that of the baselines.
-We implement various GPU kernels based on the algorithms in Section 4.4, and we present the best number among all implementations for each message size as we do for NCCL, RCCL, and MSCCL as well.
-All NCCL, RCCL, and MSCCL numbers are fine-tuned for each environment and message size by adjusting their environment variables, such as the number of channels (affects the number of threads), chunk size (affects the size of data to be transferred at once), type of algorithm (such as ring, tree, or NVLS [23]), the topology (XML file that describes the intra-node link topology of GPUs), etc.
-For MSCCL, we use the fastest algorithm for each buffer size [20].
-We leverage NCCL’s user buffer registration (i.e., ncclMemAlloc[26] and the CUDA/HIP Graph features for best performance.
+5.1. Collective Communication
+To compare the performance of collective communication, we implement the AllReduce algorithms described in Section 4.4 using the MSCCL++ DSL API.
+We also implement various AllGather algorithms in a similar manner, and we omit the implementation details.
+We present the best number among all implementations for each message size and environment.
+All NCCL, RCCL, and MSCCL numbers are fine-tuned for each environment and message size by adjusting their environment variables, such as the number of channels (affects the number of threads), chunk size (affects the size of data to be transferred at once), type of algorithm (such as ring, tree, or NVLS (NVIDIA, 2025c)), the topology (XML file that describes the intra-node link topology of GPUs), etc.
+For MSCCL, we use the fastest algorithm for each message size (Microsoft, 2025b).
+We use NCCL’s user buffer registration API (i.e., ncclMemAlloc(NVIDIA, 2025f) and the CUDA/HIP Graph APIs for best performance.
 For visibility, we separate the range of message sizes into small (up to 1MB, presented latency) and large (1MB and above, presented algorithm bandwidth (AlgoBW) [^6] Defined as the message size divided by the latency.) in the figures.
 The small message sizes represent inference scenarios (such as LLM token sampling a.k.a.
-decode [30]), while the large message sizes represent both training and inference scenarios (such as gradient accumulation during back-propagation, or LLM prompt processing, a.k.a.
-prefill [30]).
-Figure 8: AllReduce, A100-40G, single-node (8 GPUs).
-Figure 9: AllReduce, A100-40G, 2-node (16 GPUs).
-Figure 10: AllReduce, A100-40G, 4-node (32 GPUs).
-Figure 11: AllReduce, H100, single-node (8 GPUs).
-Figure 12: AllReduce, MI300x, single-node (8 GPUs).
-We compare the AllReduce performance over A100-40G nodes in Figure 8, 9, and 10.
-All data points are validated multiple times with separate trials.
-MSCCL++ outperforms baselines for both small messages (up to 3.5x and 2.1x faster over NCCL and MSCCL, respectively) and large messages (up to 1.6x and 1.4x faster over NCCL and MSCCL, respectively).
-MSCCL is faster than NCCL in most cases, and all the benefit comes from better algorithms.
-The gain is especially big for small message sizes where MSCCL uses all-pairs algorithms, while NCCL uses the ring algorithm that is worse in terms of latency.
-The gap between MSCCL and NCCL shrinks for very large message sizes where the performance is limited by the link bandwidth.
-MSCCL++ provides further benefits over MSCCL by implementing the algorithms in a more efficient way via the primitive API.
-All MSCCL and MSCCL++ numbers in the figures use the same collective algorithms and differ only by implementation in all cases.
-Even if MSCCL and MSCCL++ separately select the best-performing algorithm for each case, they come up with the same algorithm.
-Therefore, the gap between MSCCL and MSCCL++ directly shows the benefits of the proposed primitive API.
+decode (Pope et al., 2023)), while large message sizes represent both training and inference scenarios (such as gradient accumulation during back-propagation, or LLM prompt processing a.k.a.
+prefill (Pope et al., 2023)).
+Results.
+Figure 8 compares the AllReduce performance over A100-40G.
+MSCCL++ outperforms baselines for both small messages (up to 4.2x and 3.1x faster over NCCL and MSCCL, respectively) and large messages (up to 1.8x faster over both NCCL and MSCCL).
+Similarly, Figure 9 compares the AllGather performance over A100-40G.
+For small messages, MSCCL++ outperforms NCCL and MSCCL by up to 5.4x and 2.3x, respectively, and for large messages, MSCCL++ outperforms NCCL and MSCCL by up to 1.8x and 1.4x, respectively.
+For AllGather, we observe a few cases where MSCCL slightly outperforms MSCCL++ by up to 8.0%, which is introduced by a performance bug of the DSL API, not by our algorithm implementation.
+For AllReduce, MSCCL++ is always the fastest.
+Figure 8. AllReduce, A100-40G. MnNg means M nodes and N GPUs. X-axis is message size (at the bottom of the figure).
+Gain Breakdown.
+The performance differences can come from both high-level differences in algorithms (i.e., design of data flow routes and pipelines) and low-level optimizations (i.e., how to transfer data, synchronize, and manage buffers).
+The former benefit is already delivered by MSCCL that enables the customization of algorithms over the NCCL communication stack.
+Therefore, the key benefits of MSCCL++ lie in the latter by exposing the primitive functionalities to users.
+Since MSCCL and MSCCL++ use the same high-level algorithm for all cases in Figure 8 and 9, the performance gap between the two comes purely from the benefits of MSCCL++.
+NCCL vs MSCCL.
+MSCCL is faster than NCCL in most cases, and all the benefit comes from better algorithms customized for our environments.
+The gain is especially big for small message sizes, where MSCCL uses an all-pairs algorithm (which is not supported by NCCL), while NCCL uses the ring algorithm that is worse in terms of latency.
+For very large message sizes $\geq$ 256MB in multi-node cases, MSCCL achieves higher bandwidth than NCCL because it uses a hierarchical algorithm (2PH), which is more bandwidth-efficient.
+Figure 9. AllGather, A100-40G. MnNg means M nodes and N GPUs. X-axis is message size (at the bottom of the figure).
 […]
-Figure 11 and 12 compare the single-node AllReduce performance over H100 and MI300x nodes.
-Similarly, MSCCL++ outperforms baselines by up to 3.8x and 2.2x for small and large message ranges, respectively.
-This shows that MSCCL++ is effective across different GPU architectures and links.
-The speedup over NCCL/MSCCL for large message sizes on H100 is especially interesting, because the algorithm used by MSCCL++ here is 2PAM (Section 4.4), which simply calls two NVLS multimem instructions element-wise (one for reduction and one for broadcast) in a loop.
-This shows that NCCL primitives are built with a big overhead of unnecessary components that can be avoided in MSCCL++.
-5.2 LLM Inference Acceleration
-Figure 13: Speedup for decodes with tensor parallelism = 8
+DSL vs Primitive.
+From our 9-month experience using the MSCCL++ DSL API, we find that it reduces development time from weeks to days, compared to using the Primitive API directly.
+The DSL helps automate initialization, check for mistakes, and provide performance analysis.
+We have implemented both the DSL and Primitive versions up to 2-node experiments to verify that their performance is similar, while 4-node experiments are done only using the DSL.
+However, as the DSL introduces higher level instructions than the Primitive API, DSL versions perform 3% worse than the Primitive versions on average, and are up to 18% worse in one corner case.
+Despite limitations, MSCCL++ DSL is still useful for quick prototyping of collective communication and easy-to-understand algorithm description.
+Collective vs Primitive. For AI inference workloads, we implement the best algorithms in our collective kernels using the SpeCL Primitive API, so there is no performance difference.
+Figure 10. Speedup for decodes with tensor parallelism = 8
+5.2. LLM Inference
 […]
-6 Related Work
-NVIDIA Collective Communications Library (NCCL) [21] and AMD RoCm Collective Communications Library (RCCL) [4] are vendor supplied libraries for NVIDIA and AMD GPUs respectively.
+5.3. Cross-Hardware Support
+To show the effectiveness of MSCCL++ on different hardware platforms, Figure 11 and 12 compare the single-node AllReduce performance on H100 and MI300x.
+[…]
+Figure 11. AllReduce, H100, single-node (8 GPUs).
+Figure 12. AllReduce, MI300x, single-node (8 GPUs).
+MI300x.
+MSCCL++ outperforms RCCL and MSCCL by up to 3.8x and 1.9x for small messages, and by up to 2.2x and 1.6x for large messages, respectively.
+Our algorithm implementations take special considerations to best utilize the Infinity Fabric that peer-to-peer connects all GPUs in a node, unlike NVLink that connects all GPUs to a switch.
+This means, for best link utilization, we need to copy data to all peers at the same time as much as possible, unlike for NVIDIA GPUs where we can copy data to each peer one-by-one back-to-back.
+Using the MSCCL++ DSL, this is as easy as changing the order of two nested for loops, one that walks through data elements and the other that walks through different peers.
+6. Related Work
+NVIDIA Collective Communications Library (NCCL) (NVIDIA, 2025a) and AMD ROCm Collective Communications Library (AMD, 2025) (RCCL) are vendor supplied libraries for NVIDIA and AMD GPUs respectively.
 Cowan et al.
-[9] improved over these libraries by allowing execution of collective algorithms specialized for a size and topology.
+(Cowan et al., 2022) improved over these libraries by allowing execution of collective algorithms specialized for a size and topology.
 In this paper, we showed that primitives provided by MSCCL++ performs better than the primitives of these libraries.
 The primitive interface of MSCCL++ has similarities with some other works in terms of being a GPU-side API for communication.
-For example, the NVIDIA OpenSHMEM (NVSHMEM) [22] is a parallel work that provides primitive functions such as nvshmemx_putmem_warp as well as a few collective communication functions such as nvshmemx_broadcastmem_warp.
-However, we could not find any implementation where NVSHMEM outperforms NCCL (or MSCCL++) for collective communication, and qualitative comparison is difficult as NVSHMEM is not open source.
-ARK [13] is another work that proposes a GPU-side control plane for communication, but it is implemented as a monolithic end-to-end ML system rather than a standalone communication library.
+For example, the NVIDIA OpenSHMEM (NVSHMEM) (NVIDIA, 2025b) is a parallel work that provides primitive functions such as nvshmemx_putmem_warp as well as a few collective functions such as nvshmemx_broadcastmem_warp.
+However, we could not find any implementation where NVSHMEM outperforms NCCL (or MSCCL++) for collective communication, and quantitative comparison is difficult as NVSHMEM is not open source.
+ARK (Hwang et al., 2023b) is another work that proposes a GPU-side control plane for communication, but it is implemented as a monolithic end-to-end ML system rather than a standalone communication library.
 UCX, an open-source framework developed by Shamis et al.
-[33] and libraries implementing the MPI standard [35] offer a comprehensive set of APIs for high-performance computing (HPC) environments, however, MSCCL++ is more specific to GPU acceleration.
-Works like SCCL [6], TACCL [32], and TE-CCL [18] aim to accelerate GPU collective communication by synthesizing efficient data transfer algorithms for performing collective communication, while still using the vendor-provided primitives underneath.
-The primitives interface provided by MSCCL++ can be used with the generated algorithms to capture their benefit.
-Finally, LLM inference frameworks like vLLM [16] and TensorRT-LLM [25] implement custom AllReduce kernels for achieving high performance communication.
+(Shamis et al., 2015) and libraries implementing the MPI standard (Thakur et al., 2005) offer a comprehensive set of APIs for high-performance computing (HPC) environments.
+In contrast, MSCCL++ is more specific to GPU acceleration.
+SCCL (Cai et al., 2021), TACCL (Shah et al., 2023), and TE-CCL (Liu et al., 2024) aim to accelerate GPU collective communication by synthesizing efficient data transfer algorithms for performing collective communication, while still using the vendor-provided primitives underneath.
+The MSCCL++ primitives can be used with the generated algorithms to capture their benefit.
+Finally, LLM inference frameworks like vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2025e) implement custom AllReduce kernels for achieving high performance communication.
 However, the custom implementations are not general-purpose and often only limited to single-node collective communication.
 MSCCL++, on the other hand, enables communication across multiple types of interconnects and has equivalent AllReduce performance to vLLM’s custom single-node AllReduce.
-7 Conclusion
+7. Conclusion
 MSCCL++ is a novel GPU communication stack designed for high-performance AI applications.
@@
 Moreover, such a design can reduce the overall development and optimization effort for GPU communication, and accelerates adoption of fast evolving hardware technologies.
-By implementing collective communication using the proposed MSCCL++ interfaces, we can achieve up to 3.8×3.8$\times$3.8 × speedup for standalone collectives, and up to $15\%$ speedup for end-to-end AI inference.
+By implementing collective communication using the proposed MSCCL++ interfaces, we can achieve up to 5.4×5.4$\times$5.4 × speedup for standalone collectives, and up to $15\%$ speedup for end-to-end AI inference.